### v3

# Channel Measurement, Modeling, and Simulation for 6G: A Survey and Tutorial

Abstract
The sixth generation (6G) mobile communications have attracted substantial attention in the global research community of information and communication technologies (ICT). 6G systems are expected to support not only extended 5G usage scenarios, but also new usage scenarios, such as integrated sensing and communication (ISAC), integrated artificial intelligence (AI) and communication, and communication and ubiquitous connectivity. To realize this goal, channel characteristics must be comprehensively studied and properly exploited, so as to promote the design, standardization, and optimization of 6G systems. In this paper, we first summarize the requirements and challenges in 6G channel research. Our focus is on channels for five promising technologies enabling 6G, including terahertz (THz), extreme MIMO (E-MIMO), ISAC, reconfigurable intelligent surface (RIS), and space-air-ground integrated network (SAGIN). Then, a survey of the progress of the 6G channel research regarding the above five promising technologies is presented in terms of the latest measurement campaigns, new characteristics, modeling methods, and research prospects. Moreover, a tutorial on the 6G channel simulations is presented. We introduce the BUPTCMCCCMG-IMT2030, a 6G link-level channel simulator, developed based on the ITU/3GPP 3D geometry-based stochastic model (GBSM) methodology. The simulator supports the channel simulation of the aforementioned 6G potential technologies. To facilitate the use of the simulator, the tutorial encompasses the design framework, user guidelines, and application examples. This paper offers in-depth, hands-on insights into the best practices of channel measurements, modeling, and simulations for the evaluation of 6G technologies, the development of 6G standards, and the implementation and optimization of 6G systems.
Index Terms:
6G, channel modeling, channel measurement, channel simulation, terahertz, MIMO, integrated sensing and communication, reconfigurable intelligent surface, space-air-ground integrated network
I Introduction
I-A Overview of 6G
As the fifth generation (5G) mobile communication networks have been deployed worldwide, studies on the sixth generation (6G) mobile communication technologies, which are expected to be rolled out for 2030 and beyond, are being vibrantly conducted. Several 6G research projects and organizations have been established. Representative examples include the Finnish “6Genesis Flagship Program” [1], the European “Hexa-X” project [2], the Chinese “IMT-2030 (6G) Promotion Group” [3], and the North American “Next G Alliance” [4]. Simultaneously, international standardization organizations have launched several key initiatives. For instance, in 2022, the International Telecommunication Union (ITU) identified future technology trends [5] for International Mobile Telecommunications (IMT) systems toward 2030 and beyond. Very recently, the recommendation framework for IMT-2030, which is commonly known as the 6G vision, was finalized [6]. This means that the global 6G standardization campaign has been formally initiated. With the recommendation framework, ITU establishes the overall objectives, procedures and timelines for the development of 6G mobile communication standards. Following the ITU requirements and guidelines, the third Generation Partnership Project (3GPP) is expected to start the drafting of the 6G specifications in approximately 2025, to prepare a technology proposal for the ITU radiocommunication sector (ITU-R) afterwards, and to support the completion of 6G standardization no later than 2030.
6G aims to establish a more inclusive information society than 5G. Its goal is to intelligently connect humans, machines, and various other things [6]. Some key drivers of 6G have been discussed in [5]. These drivers can be understood from the perspective associated with continuously growing mobile traffic and emerging applications. Specifically, it is forecast that by 2030, global mobile traffic will reach 5016 exabytes (EB) per month [7], which means a growth of approximately 80 times compared with that of 2020. The potential emerging applications of 6G [5] can be categorized into several main usage scenarios [6]. One extends enhanced mobile broadband (eMBB), which was introduced in 5G, to target immersive applications, such as holographic communications, tactile and haptic internet applications, and extended reality (XR). Similarly, hyperreliable and low-latency communications and massive communications are extensions of 5G usage scenarios of ultrareliable and low latency communications (URLLC) and massive machine type communications (mMTC), respectively. In addition, three other usage scenarios have been identified, i.e., ubiquitous connectivity, integrated artificial intelligence (AI) and communication, and integrated sensing and communication (ISAC) [6]. The applications of integrated AI/sensing and communication include perceptive mobile networks, digital twins, and the proliferation of networked AI. As a result, the capabilities of 6G are enhanced across multiple dimensions, including data rate, spectrum efficiency, reliability, latency, mobility, and connection density [8, 9, 10, 11, 12, 13, 14, 6]. Moreover, new capabilities are considered, such as all-domain coverage, sensing-related and AI-related capabilities, sustainability and computing power [6].
I-B Requirements and Challenges of 6G Channel Research
The radio channel channel is the transmission medium between a transmitter (Tx) and receiver (Rx). It plays a crucial role in determining the ultimate performance limit of mobile communications systems [15, 16, 17]. As shown in Fig. 1, 6G communication systems will incorporate various novel technological components [5, 16, 18, 10, 19, 20, 21, 8, 22], such as the extension of the frequency band to terahertz (THz) [23, 24, 25, 26, 17, 27], the extension of the antenna array to the extreme MIMO (E-MIMO) [9, 22, 28, 29], the integration of sensing into the communication system (i.e. ISAC) [30, 31, 32], the introduction of a programmable wireless environment by using reconfigurable intelligent surface (RIS) [33, 34, 35], and the extension of coverage to space-air-ground integrated network (SAGIN) [36, 37, 38]. Each of these may result in distinct requirements and characteristics for the wireless channel, as summarized below.
Figure 1: Illustration of 6G channel research requirements.
I-B1 Higher frequency band, propagation attenuation, and channel sparsity
The deployment of 5G has already achieved highly efficient utilization of the sub-6 GHz frequency band and harnessed millimeter wave (mmWave) frequencies to deliver mobile devices with unparalleled multi-Gbps data rates [39]. The THz frequency band (0.3-3 THz according to ITU radio regulations or 0.1-10 THz as a commonly used wider range) possesses an atmospheric transmission window with low molecular absorption losses, and abundant bandwidth having been allocated to terrestrial mobile communication services at the World Radiocommunication Conference (WRC) in 2019 [40]. As the era of 6G begins, the demand for a peak data rate of 1 Tbps requires a comprehensive exploitation of the THz frequency band [16]. Prototype demonstration and testing systems for THz communication have already been developed to achieve high-speed transmission. For instance, high-capacity wireless transmission at 1.44 Tbps has been achieved within the THz frequency band of 135-170 GHz [41]. However, it is essential to recognize that the THz band has distinctive characteristics in contrast to the sub-6 GHz and mmWave frequency bands. The shorter wavelength ($\leq$ 3 mm) signals in the THz band experience the roughness of the objects in the environment. Previously smooth surfaces, such as walls and tree leaves, appear rougher at THz band. The specular reflection may lose its privileged position, allowing diffuse scattering to become the predominant reflection behavior. In general, the THz wave experiences severe transmission losses and molecular absorption, thereby causing changes in the path loss characteristics and the emergence of sparsity. Sparsity refers to the condition where only a small number of multipath components (MPCs) carry the majority of the received signal power. Currently, the ITU/3GPP channel model [42, 43] only supports frequencies up to 100 GHz for 5G. Therefore, it is urgent to establish a comprehensive THz channel model, which is capable of covering the typical scenarios (such as indoor hotspot (InH) and urban microcell (UMi)) and incorporating the distinctive THz channel characteristics (such as high path loss and sparsity).
I-B2 Larger-scale antenna array, near-field effect, and spatial non-stationarity
It is the 1970’s that researchers suggested to introduce MIMO technologies into wireless communication systems. In 1995, Telatar theoretically derived the MIMO channel capacity under the fading condition [44]. It was demonstrated that the MIMO technology can substantially increase the capacity of wireless communication system. Then, the MIMO technology was introduced into the fourth generation (4G) mobile communication system, which can support eight array elements and highly improve the system spectrum efficiency [45]. However, the channel model used in 4G systems only supports the azimuth dimension. With the proposal of massive MIMO, three-dimensional (3D) MIMO [46] is a typical massive antenna realization. Therefore, the elevation dimension was introduced into the ITU/3GPP 5G standard channel model[42, 43]. The 5G channel model captures the elevation-azimuth-time-frequency channel characteristics, and meets the requirements of 3D MIMO and massive MIMO technologies. In the 6G era, the number of array elements is expected to increase significantly, reaching a range from a few hundreds to thousands. As a result, the E-MIMO technology based on the much larger-scale antenna arrays is envisaged to be deployed. The E-MIMO with 1024$\times$1024 array elements can offer abundant spatial degrees of freedom for supporting 100 times peak data rate improvement in 6G mobile communication systems [47]. However, some new channel characteristics emerge in E-MIMO systems. To elaborate a little further, when the propagation distance is shorter than the far-field Rayleigh distance, the plane wave assumption is invalid in the near-field region [48, 49]. Therefore, the the near-field spherical wave and spatial non-stationary (SnS) characteristics [26, 50] cannot be ignored in E-MIMO communication systems. As shown in Fig. 1, the E-MIMO can be modeled using near-field spherical propagation in the near-field region. In addition, with the expansion of physical array aperture, different array elements see different scatterers, resulting in SnS on the array. However, these characteristics are not covered by the ITU/3GPP 5G channel model framework. Therefore, it is necessary to investigate the impact of the near-field spherical wave and the SnS characteristic through channel measurements and simulations for 6G channel modeling.
I-B3 ISAC, simultaneous presence of sensing channel and communication channel
ISAC is one of the candidate technologies for 6G[51]. It is capable of exploiting the potential sensing ability of base stations (BSs) to obtain environmental information to assist communication, thus benefiting beam alignment, directional coverage, etc. At the same time, multipath information in communication signals can also enhance sensing ability. In ISAC systems, sensing aims to obtain the state information, such as the speed and location, of the targets in the particular environment through echoes and the collaboration between BSs [52]. Therefore, in the study of ISAC channels, it is essential to introduce the radar cross section (RCS), a target-related feature, and consider the impact of the environment on the sensing channel [53]. Moreover, due to the simultaneous use of the propagation environment by the communication and sensing signals, some shared propagation scatterers may be experienced in the ISAC system [53, 54]. The clusters contributed by these shared scatterers may coexist and disappear in both communication and sensing channels. Hence, the communication channel component and the sensing channel component of the ISAC channel may be partially related to each other. These novel propagation characteristics facing the ISAC technology cannot be adequately captured by existing communication channel models. Therefore, an ISAC channel model incorporating these characteristics based on realistic channel measurements is of great significance for the design and performance evaluation of ISAC systems.
I-B4 RIS, reflection coefficient characteristics, and cascaded channel
RIS is considered as a potential technology that can overcome the challenges in traditional beam-space antenna array based beamforming [6]. It is capable of significantly improving the channel gain by actively controlling the wireless propagation environment [14]. For example, when the line-of-sight (LOS) path is obstructed, RIS can provide a high-power reflection path. Additionally, it promises low cost and high energy efficiency. However, since RIS can program the radio propagation environment to some degree, the role of the RIS itself should be taken into account when studying the wireless channels that involve RIS. This makes the traditional wireless channel model no longer suitable for the cascaded channel of RIS, and the channel modeling theory for RIS is not yet mature. Furthermore, for RIS, both the decoupling of the channel model and the construction of a high-precision electromagnetic response model are urgent problems to be solved [14]. In addition, the high path loss of the RIS cascaded channel and the flexibility of RIS deployment locations pose new challenges for RIS channel measurement.
I-B5 SAGIN, meteorological effects, and dynamic properties
SAGIN is a heterogeneous architecture that integrates different networks like space networks, aerial networks, and terrestrial communication networks. It is expected to be realized in the future 6G era [55]. The terrestrial networks achieve normal coverage of various scenarios in urban and rural areas as discussed above. Space and aerial networks, in which devices are located on the satellite, high altitude platform stations (HAPS), and unmanned aerial vehicles (UAV), can realize ubiquitous coverage in remote areas, sea, and airspace. Due to the ability to achieve all-domain coverage, SAGIN has attracted much attention from both academia and the industry. In recent years, several organizations have started their projects to deploy satellite constellations in low earth orbit (LEO). These constellations contain hundreds or thousands of satellites and are deployed to provide worldwide coverage. Examples of these constellations include the SpaceX Starlink, which is announced to comprise over 10000 satellites, the Norwegian STEAM network, which is claimed to contain 4257 satellites, and the OneWeb constellation, which is a smaller constellation with 720 satellites [38]. The ubiquitous coverage of SAGIN benefits a variety of new services and applications. It also leads to more diverse scenarios and more complex channel characteristics. Since SAGIN involves space, air, and ground, different types of communication segments should be considered, such as space-to-ground (SG), air-to-ground (AG), and others. As for the SG segment, the transmission signal may suffer from an extra loss caused by meteorological environmental factors like atmosphere, rain, clouds, and fog. When the signal travels through the ionosphere, fluctuations of amplitude and phase (scintillation), rotations of polarization angle (Faraday rotation), delay dispersion, and others are found due to the ionosphere effects. As for the AG segment, the high change rate of channel characteristics because of the UAV velocities, buildings, and ground surface composition is exhibited. Besides, characteristics of AG channels like Doppler spread should be considered. For terrestrial networks, channel characteristics are discussed above. For the channel of SAGIN, both each segment channel and the integration effect on channel characteristics require consideration and being well addressed.
I-C Paper Contributions and Organization
Research on wireless channels has a vital role in the design, standardization and optimization of 6G systems. At the time of writing, a few surveys on some aspects of 6G channels [9, 16, 24, 56, 57] have been published. Nevertheless, there has not been a comprehensive survey on the entire portfolio of 6G channel research. We not only provide a comprehensive survey of the channel models of the 6G potential enabling technologies including THz, E-MIMO, ISAC, RIS and SAGIN, but also provide a tutorial on the channel simulator BUPTCMCCCMG-IMT2030, which support for various scenario and features of the above technologies. Taking into account mainstream models, the simulator uses the 5G standard channel modeling framework ITU-R M.2412 [42] (principles described in Section II) as the basis for expansion to 6G. Our major contributions via this paper are as follows:
  • The representative THz channel measurement campaigns in the 100-1100 GHz frequency band are reviewed and analyzed. Then, a detailed discussion on the THz channel characteristics is given, specifically focusing on reflection, high loss and sparsity. Finally, the pros and cons of stochastic, deterministic, and hybrid modeling approaches for THz channels are further discussed.
  • The E-MIMO channel measurement platforms are systematically introduced, including frequency domain and time domain platforms. Then, the near-field spherical wave model of E-MIMO and the characteristics of SnS are discussed in detail. At the same time, the division of near-field range is analyzed. In addition, several channel modeling methods such as the statistical, deterministic and hybrid channel modeling are compared, then a realistic yet low-complexity SnS channel modeling framework is discussed. Finally, the future research directions of E-MIMO are pointed out.
  • A summary of research progress on ISAC channels in both monostatic and bistatic sensing modes is presented. Furthermore, an in-depth analysis of the novel characteristics introduced by the ISAC channel is conducted. To capture the channel sharing feature, we review a cluster-based ISAC channel model. Finally, the pros and cons of various modeling approaches are discussed for ISAC channels.
  • Some measurement campaigns related to the RIS channel are summarized, especially regarding the power gain brought by RIS. In addition, the characteristics of the RIS channel are introduced in detail, including the response characteristics of the important intermediate node RIS and the new loss characteristics presented due to the cascading of two subchannels and RIS response. Then, to accurately describe the RIS channel, we provide various types of RIS channel modeling approaches, including statistical, deterministic, and hybrid approaches. Finally, future research directions are discussed.
  • Channel measurements in SAGIN are reviewed, including the SG, AG, and air-to-air (AA) channels, which is followed by a detailed description of the channel characteristics of SAGIN. In addition, channel models for simulating SAGIN are introduced. Finally, a brief summary of potential research directions for future work is presented.
  • The 6G channel simulator BUPTCMCCCMG-IMT2030 developed by our team is introduced. The design framework and user guidelines of the simulator are detailed. We also provide simulation examples to characterize THz, E-MIMO, ISAC, RIS, and SAGIN.
Figure 2: Organization of the paper.
The rest of this paper is organized as follows. In Section III–VII, we provide an overview of the state-of-the-art channel research on THz, E-MIMO, ISAC, RIS and SAGIN technologies in terms of measurements, characteristics and modeling and discuss the prospects. Section VIII presents an tutorial on the 6G channel simulator BUPTCMCCCMG-IMT2030 developed by our team. Finally, conclusions are drawn in Section IX. For better clarity, the organization of the paper is shown in Fig. 2.
II Modeling Methodology of ITU/3GPP Stochastic Channel
The transmitted signals through the wireless channel will suffer propagation losses, which can be roughly divided into two types, large-scale and small-scale channel fading [58]. In practice, both types of fading occur [15] and can significantly affect all aspects of the communication system [16]. Large-scale fading affects the received signal strength over long Tx-Rx distances (typically hundreds or thousands of meters). Hence, it impacts coverage prediction and interference analysis, and it is an essential factor for wireless communication system deployment. However, small-scale fading generally refers to the fluctuation of signal intensity for short distances of a few wavelengths or short periods of seconds [59]. Small-scale fading is more relevant to the design of reliable and efficient communication systems [58]. The characteristics, such as the delay and the angle of MPCs, play a vital role in designing physical-layer transmission techniques [60]. In this section, we briefly introduce the modeling principles of the 5G stochastic standard channel model (ITU-R M.2412 [42] and 3GPP TR 38.901 [43]) from large-scale fading and small-scale channel fading perspectives.
II-A Large-Scale Channel Fading
Large-scale fading comes from path loss as a function of distance and shadowing by large objects such as buildings and hills [58]. The path loss represents the reduction in radio wave power over a long distance [59]. Shadow fading is mainly due to the multipath blockage, which also causes significant attenuation of the signal strength. The 5G standard provides three types of path loss models, including the close-in (CI) model, floating-intercept (FI) model, and alpha-beta-gamma (ABG) model [61, 62, 63]. The FI model is expressed as
$\textrm{PL}_{\textrm{FI}}(d)[\textrm{dB}]=\beta_{\textrm{FI}}+10\alpha_{\textrm{FI}}\textrm{lg}d+X_{\sigma}^{\textrm{FI}},$ (1)
where $\textrm{PL}_{\textrm{FI}}(d)$ denotes the path loss in dB, which is a function of the Tx-Rx separation distance $d$. $\beta_{\textrm{FI}}$ and $\alpha_{\textrm{FI}}$ are two fitted parameters. $X_{\sigma}^{\textrm{FI}}$ is a zero-mean Gaussian variable with standard deviation $\sigma$ in dB representing shadow fading. Note that $\beta_{\textrm{FI}}$ is frequency dependent.
The CI path loss model can be provided by (2).
$\begin{split}\textrm{PL}_{\textrm{CI}}(f,d)[\textrm{dB}]=&20\textrm{lg}\frac{4\pi d_{0}f}{\textrm{c}}+10\alpha_{\textrm{CI}}\ \textrm{lg}\frac{d}{d_{0}}+X_{\sigma}^{\textrm{CI}}\\ &\textrm{for}\ d\geq d_{0},\ d_{0}=1\,\textrm{m},\end{split}$ (2)
where $f$ is the center frequency, $d$ is the Tx-Rx separation distance greater than the reference distance $d_{0}$, c denotes the speed of light, $\alpha_{\textrm{CI}}$ is the single model parameter known as the path loss exponent (PLE), and $X_{\sigma}^{\textrm{CI}}$ is the zero-mean Gaussian variable, which describes the large-scale signal fluctuations about the mean path loss over distance.
The third type of path loss model, ABG model is represented as a function of frequency and distance [62, 64]. It can describe the path loss at various frequencies, and is expressed as
$\begin{split}\textrm{PL}_{\textrm{ABG}}(f,d)[\textrm{dB}]=&10\alpha_{\textrm{ABG}}\ \textrm{lg}\frac{d}{d_{0}}+\\ &\beta_{\textrm{ABG}}+10\gamma_{\textrm{ABG}}\ \textrm{lg}\frac{f}{1\ \textrm{GHz}}+X_{\sigma}^{\textrm{ABG}}\\ &\textrm{where}\ \ d_{0}=1\,\textrm{m},\end{split}$ (3)
where $d_{0}$ is the received power reference point, typically set to 1 m, here $f$ is in GHz. $\alpha_{\textrm{ABG}}$, $\beta_{\textrm{ABG}}$ and $\gamma_{\textrm{ABG}}$ are fitting parameters, where $\alpha_{\textrm{ABG}}$ and $\gamma_{\textrm{ABG}}$ characterize the distance dependence and frequency dependence of path loss, respectively, and $\beta_{\textrm{ABG}}$ is an offset value in dB. The typical value of $\gamma_{\textrm{ABG}}$ is 2, which corresponds to free space propagation. $X_{\sigma}^{\textrm{ABG}}$ is the zero-mean Gaussian variable. In the ABG model, the path loss varies with the frequency and distance. In [65], it was demonstrated that the CI and ABG models offer comparable modeling performance upon using field measurement data.
II-B Small-Scale Channel Fading
Geometry-based statistical channel modeling is the mainstream approach adopted by both ITU and 3GPP in 4G and 5G. According to the definition of the geometry-based stochastic model (GBSM) [66], the channels are generated by summing the contributions of rays based on geometry principles and specific small-scale statistical channel parameters from measurement. Geometry-based modeling enables the channel to be independent of the antenna and allows the configuration of different antennas and different element patterns. Additionally, for individual channel snapshots, the channel parameters, such as delay, power, angle of arrival, and angle of departure, are stochastically determined based on statistical distributions extracted from channel measurements. In the GBSM-based channel model of the 5G standard, a new dimension of elevation are introduced at both Tx and Rx, enabling support for 3D MIMO.
The 3D GBSM based channel model is illustrated in Fig. 3. The small sphere with several dots inside represents a scattering region that generates a cluster of propagation paths. Each cluster is constituted by $M$ rays, and $N$ clusters are assumed. Tx and Rx are equipped with $S$ and $U$ antennas, respectively. The small-scale parameters (SSPs), such as the delay $\tau_{n,m}$, zenith angle of arrival $\theta_{n,m,\textrm{ZOA}}$, azimuth angle of arrival (AOA) $\phi_{n,m,\textrm{AOA}}$, azimuth angle of departure (AOD) $\phi_{n,m,\textrm{AOD}}$, and zenith angle of departure (ZOD) $\theta_{n,m,\textrm{ZOD}}$, are assumed to be different for the $m$th ray in the $n$th cluster. Radio waves propagate in 3D space, and MPCs exhibit not only horizontal-dimension characteristics but also vertical-dimension characteristics.
Figure 3: Illustration of the 3D GBSM-based channel model [67].
In the ITU-R 5G stochastic channel model, the channel impulse response (CIR) for the ray from the $s$th transmitting antenna to the $u$th receiving antenna of the $n$th cluster is a function of time $t$, delay $\tau$, angle of arrival $\bm{\Omega}$ and angle of departure $\bm{\Psi}$ given by
$\begin{split}&h_{u,s,n}(t,\tau,\bm{\Omega},\bm{\Psi})=\sum_{m=1}^{M_{n}}a_{n,m}F_{\textrm{Rx},u}(\bm{\Omega}_{n,m})\\ &\textrm{exp}(j2\pi\lambda^{-1}(\hat{r}_{\textrm{Rx},n,m}\cdot\vec{d}_{\textrm{Rx},u}))F_{\textrm{Tx},s}(\bm{\Psi}_{n,m})\\ &\textrm{exp}(j2\pi\lambda^{-1}(\hat{r}_{\textrm{Tx},n,m}\cdot\vec{d}_{\textrm{Tx},s}))\delta(\tau-\tau_{n,m})\\ &\delta(\bm{\Omega}-\bm{\Omega}_{n,m})\delta(\bm{\Psi}-\bm{\Psi}_{n,m}),\end{split}$ (4)
where $a_{n,m}$ is the complex amplitude of the $m$th MPC within the $n$th cluster, $F_{\textrm{Rx},u}$ and $F_{\textrm{Tx},s}$ are the $u$th antenna pattern at the Rx side and the $s$th antenna pattern at the Tx side [68, 69], respectively. $\hat{r}_{\textrm{Rx},n,m}$ is the unit vector in the spherical coordinate system at the Rx side, $\vec{d}_{\textrm{Rx},u}$ is the location vector of the $u$th antenna at the Rx side, and the subscript Tx similarly represents the Tx side. $\delta(\cdot)$ is the Dirac function, $\lambda$ is the wavelength, and $M_{n}$ is the number of delay bins. $\bm{\Omega_{n,m}}$ and $\bm{\Psi_{n,m}}$ are the angles of arrival and departure of the MPC, respectively. They include both the elevation and azimuth angles, $\theta$ and $\phi$, for 3D MIMO systems as follows:
$\begin{split}&\bm{\Omega_{n,m}}=(\theta_{n,m,\textrm{ZOA}},\phi_{n,m,\textrm{AOA}}),\\ &\bm{\Psi_{n,m}}=(\theta_{n,m,\textrm{ZOD}},\phi_{n,m,\textrm{AOD}}).\end{split}$ (5)
In practical 5G systems, antennas can be designed and configured to exhibit specific radiation patterns that enhance the system performance, and slant polarized antennas are usually used. Thus, the radiation patterns need be modeled when generating the CIR. Horizontal and vertical linear polarizations and the $+45^{\circ}/-45^{\circ}$ slant polarization are examples of orthogonal polarizations [70]. When dual polarization is utilized, (4) is changed to (6).
$\begin{split}H_{u,s,n}(t,\tau)=\sqrt{\frac{P_{n}}{M}}\sum_{m=1}^{M}\begin{bmatrix}F_{\textrm{Rx},u,\theta}(\theta_{n,m,\textrm{ZOA}},\phi_{n,m,\textrm{AOA}})\\ F_{\textrm{Rx},u,\phi}(\theta_{n,m,\textrm{ZOA}},\phi_{n,m,\textrm{AOA}})\end{bmatrix}^{\textrm{T}}\begin{bmatrix}\textrm{exp}(j\Phi_{n,m}^{\theta\theta})&\sqrt{\kappa_{n,m}^{-1}}\textrm{exp}(j\Phi_{n,m}^{\theta\phi})\\ \sqrt{\kappa_{n,m}^{-1}}\textrm{exp}(j\Phi_{n,m}^{\phi\theta})&\textrm{exp}(j\Phi_{n,m}^{\phi\phi})\end{bmatrix}\\ \begin{bmatrix}F_{\textrm{Tx},s,\theta}(\theta_{n,m,\textrm{ZOD}},\phi_{n,m,\textrm{AOD}})\\ F_{\textrm{Tx},s,\phi}(\theta_{n,m,\textrm{ZOD}},\phi_{n,m,\textrm{AOD}})\end{bmatrix}\textrm{exp}(j2\pi\frac{\hat{r}_{\textrm{Rx},n,m}^{\textrm{T}}\cdot\vec{d}_{\textrm{Rx},u}}{\lambda})\textrm{exp}(j2\pi\frac{\hat{r}_{\textrm{Tx},n,m}^{\textrm{T}}\cdot\vec{d}_{\textrm{Tx},s}}{\lambda})\textrm{exp}(j2\pi v_{n,m}t)\delta(\tau-\tau_{n,m})\end{split}$ (6)
In (6), $\left\{{\Phi}_{n,m}^{\theta\theta},{\Phi}_{n,m}^{\theta\phi},{\Phi}_{n,m}^{\phi\theta},\textrm{and}\;{\Phi}_{n,m}^{\phi\phi}\right\}$ are random initial phases corresponding to four different polarization combinations ($\theta\theta$, $\theta\phi$, $\phi\theta$, and $\phi\phi$, respectively) for ray $m$ of cluster $n$. The distribution for the initial phases is uniform within ($-\pi,\pi$). ${F}_{\textrm{Rx},u,\theta}$ and ${F}_{\textrm{Rx},u,\phi}$ are the field patterns of the receiving antenna element $u$ in the direction of the spherical basis vectors, which are characterized by $\theta$ and $\phi$, respectively. Then, ${F}_{\textrm{Tx},s,\theta}$ and ${F}_{\textrm{Tx},s,\phi}$ are the field patterns of the transmitting antenna element $s$ in the direction of $\theta$ and $\phi$. $\kappa_{n,m}$ is the cross-polarization power ratio (XPR) on a linear scale. $v_{n,m}$ is the Doppler frequency component of ray $m$ of the $n$th cluster. $P_{n}$ is the normalized power of the $n$th cluster.
III THz Channel Measurements and Models
As presented in Section I, 6G uses a higher frequency spectrum than previous generations [6] to achieve multi-Gbps data rate services such as immersive communication. Recently, new frequency bands beyond 5G, such as mid bands [71] and THz bands [72], have been researched to meet the requirement of the large data rates. Among these frequency bands, THz has been identified as a promising spectrum in 6G due to the wide bandwidth and low interference characteristics, and the ITU has shown strong support for the research and development of THz technologies. As shown in Fig. 4, the ITU has allocated a total of 234.2 GHz in the 100-450 GHz frequency band for fixed services and land mobile applications. Specifically, 12 bands (total 97.2 GHz bandwidth) within the 100-275 GHz frequency range [40] have been allocated by ITU. Of note, the largest contiguous bandwidth available in the frequency bands is 23 GHz. Below 200 GHz, the largest contiguous bandwidth currently allocated for fixed or mobile applications is 12.5 GHz. The limited availability of contiguous bandwidths poses a challenge for developing high-bandwidth wireless communication systems in 6G. Therefore, by the ITU Radio Regulations 2020, the WRC 2019 (WRC-19) identified 275-296 GHz, 306-313 GHz, 318-333 GHz, and 356-450 GHz, for a total of 137 GHz, which can be used for fixed service and land mobile applications [73]. Moreover, the maximum available continuous bandwidth is 94 GHz.
The WRC is not the only organization planning for THz spectrum. Also, spectrum planning and allocation regulators in different countries and regions have released parts of the THz band to facilitate scientific research. For example, Japan’s Ministry of Information and Communication (MIC) designated the frequency range of 116-134 GHz as “Commercial Telecommunications Services” in 2015, and the 152-164 GHz and 287.5-312.5 GHz bands were made available for experimental testing in 2020 [74]. In 2018, the European Conference of Postal and Telecommunications Administrations (CEPT) and the European Radiocommunications Committee (ERC) released a recommendation on “short range devices,” also known as “unlicensed devices,” which specified provisions for these devices to operate within the frequency ranges of 122-122.25 GHz and 244-246 GHz [75]. The Federal Communications Commission (FCC) decision in March 2019 authorized the unlicensed use of 116-123 GHz, 174.8-182 GHz, 185-190 GHz, and 244-246 GHz in the United States [76]. In conclusion, a high priority has been placed on THz spectrum development and utilization in various countries and organizations.
Figure 4: Accessible bandwidths for fixed service and land mobile applications from 100 to 450 GHz.
III-A THz Channel Measurements
It is important to conduct extensive THz measurements to build channel model for THz communication [77]. Up to now, many research organizations and institutions conduct multiple THz channel measurement campaigns in the abovementioned frequency bands. We have classified and summarized some representative THz channel measurement campaigns by frequency bands, striving to comprehensively cover all currently measured THz frequency bands (100-1100 GHz), as shown in Table I.
In the THz spectrum allocated for land mobile applications, as illustrated in Fig. 4, THz measurement campaigns are relatively concentrated around the frequencies of 100, 140, 220, and 300 GHz [78, 17, 26, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125, 98, 126, 127, 128, 129, 130, 131, 132, 133, 134, 109]. The frequency of 100 GHz represents the upper limit of the ITU channel model. The primary objective of most studies is to investigate the applicability of existing channel models in the THz frequency band. Due to the non-linear factors of atmospheric absorption, the selection of actual frequencies primarily considers the “atmospheric window” frequency bands. In the THz frequency band, the “atmospheric window” refers to special frequency bands, such as 140 GHz and 220 GHz, where propagation attenuation is relatively low. Hence, these bands have been the focus of extensive research. Additionally, 300 GHz falls within the lower boundary of the THz frequency range defined by the ITU radio regulations, making it a frequency band of significant interest. However, in the 450-1100 GHz band [130, 135, 136, 137], there are fewer channel measurement campaigns limited by capabilities of the hardware, such as instruments and equipment. Due to limitations in transmitted power, the majority of studies focus only on reflection, scattering, and attenuation characteristics.
Secondly, the measurement platforms in the summarized THz channel measurement campaigns are introduced. Most measurement platforms widely used in THz channel measurement campaigns are frequency domain vector network analyzer (VNA)-based platform [138]. In the VNA-based platform, the Tx and Rx are directly connected to the VNA. The simple structure, the synchronized Tx and Rx, the large bandwidth, and the high dynamic range make this measurement platform occupy the mainstream. However, due to the high loss of the cable, the measurement distance is usually short but can be extended to 100 m by radio-over-fiber (RoF) [109]. To conduct long-distance channel measurements, correlation-based time-domain channel measurement platform is used, which operates with the strong autocorrelation of pseudo-random noise (PN) sequence. It has also been used in several THz channel measurements. Since Tx and Rx are independent of each other and do not require cable connections, the platform is preferable for measurements in outdoor environments [79]. In addition, THz time-domain spectroscopy system (THz-TDS)-based [139] channel measurement platforms are limited by the output power and are commonly used for measurements of reflection and diffraction properties of materials.
Thirdly, the scenarios of the channel campaigns are presented. THz band channel measurements focus on desktops, meeting rooms, offices, data centers, computer motherboards, and indoor factory scenarios in the indoor scenarios, and UMi, canyon, and street in the outdoor scenarios, covering vision scenarios, such as wireless local area networks (WLAN), wireless personal area networks (WPAN), data center networks (DCN), and intradevice communication (IDC) [140]. However, the wireless backhaul/fronthaul scenario has been poorly studied [141]. Most of the channel measurement campaigns in these scenarios are in the 100-450 GHz frequency bands due to the limited hardware performance in the high-frequency bands. Furthermore, studies include the channel characteristics, which will be described in detail in the next subsection.
TABLE I: Summary of main THz channel measurement campaigns
Frequency
/BW (GHz)
Measurement
platforms
Scenarios Channel characteristics Reference
100 - 450 GHz
100/1.2
Correlation-
based
Office
PDP, PL, DS
[78]
100/6
VNA+RoF Empty room SnS [26]
120/40
VNA Computer motherboard PL, DS, reflection/penetration loss [87]
135/10
Data center
PL, SF, DS, AS, cluster numbers,
intra/inter-cluster parameters
[86]
140/60
Indoor
PL, SF, DS
[80]
142/1 Correlation- based
Office, conference,
classrooms, hallways,
factory building
PL, SF, DS, AS, cluster numbers,
intra/inter-cluster parameters
[100, 101]
Building materials
Reflection, scattering,
transmission, PL
[102]
UMi
PL, foliage attenuation,
cluster numbers, DS, AS,
spatial consistency, coverage
[103, 104, 105, 106]
143.1/4
VNA+RoF
Shopping mall,
airport check-in hall
PADP, PDP, PAS, PL, DS, AS
[88, 89]
144.75/7.5
140.5/1
145.5/1
Outdoor (courtyard,
linear route,
crossroad, canyon)
PDP, PL, SF, DS, AS,
Power distribution of MPC,
Qtapnumber
[91, 92]
[94, 96]
205/8
VNA
Indoor (meeting room,
office room)
PL, SF, DS, AS,
K, cluster numbers,
intra/inter-cluster parameters
[84]
[220:10:330]/2
Correlation-
based
Desktop
PL, celluar performance,
reflection
[79, 17]
275/110
VNA+RoF
Hall
PDP, PAS, PADP
[109]
280/120
VNA
Download kiosk
DS, XPR, K,
Coherence bandwidth
[98]
300/2
Correlation-
based
UMi (courtyard),
industrial environment
PDP, PL, DS, AS
[131, 133]
300/2
VNA+RoF
Spacious hall
PDP
[26]
304.2/8
Correlation-
based
V2V
Blockage, PDP, PL, DS, AS
[122]
305/10
VNA
Desktop,
building material
PL, absorption coefficients
[121]
310/20
Desktop
PDP, PL, SF, RMS DS, TCF
[111]
Computer motherboard
PL, SF, reflection
[112]
Data center
PL, reflection
[113]
313.5/15
Indoor (hallway,
corridor)
PL, SF, DS, AS, K, cluster numbers, intra/inter-cluster parameters [117, 118]
Outdoor (campus
street, atrium)
[119, 120]
450 - 1100 GHz
450/500
THz-TDS
Building material
Diffuse scattering
[130]
625/250
VNA
Desktop
Atmospheric effects, PDP
[137]
925/350
Building material
Penetration loss
[136]
1000/200
Desktop
Ground reflections
[135]
BW: bandwidth, UMi: urban microcell, V2V: vehicle-to-vehicle, PADP: power angular delay profile,
PDP: power delay profile, PL: path loss, SF: shadow fading, K: Ricean K-factor, DS: delay spread,
AS: angular spread, PAS: power angular spectrum, XPR: cross-polarization ratio, TCF: temporal correlation function
III-B Important Channel Propagation Characteristics in THz Bands
Current THz characteristics research can be mainly categorized into two types [26]. The first focuses on the material characteristics at THz bands, which actually have an impact on the radio propagation (e.g., reflection) [142]. The second is dedicated to newly emerging channel characteristics in THz bands, such as channel sparsity.
III-B1 Reflection characteristics
Research on propagation mechanism characteristics is the key foundation and fundamental premise for THz channel characteristics research. The reflection are fundamental propagation mechanism that should be thoroughly characterized to establish possible THz wireless links due to the short wavelength introduced in Section I. Therefore, this subsection focuses on the analysis and modeling of the reflection propagation characteristics of THz waves.
In the THz band, the characteristics of reflection in the case of ideally smooth and homogeneous surfaces can be described by the well-known Fresnel reflection coefficient $\Gamma$ [143]. The Fresnel reflection coefficient expressions for vertical ($\Gamma_{\textrm{v}}$) and parallel ($\Gamma_{\textrm{h}}$) polarizations for such smooth surfaces [144, 145] are expressed as:
$\Gamma_{\textrm{v}}=\frac{{Z\cos{\theta_{\textrm{in}}}-{Z_{0}}\cos{\theta_{\textrm{tr}}}}}{{Z\cos{\theta_{\textrm{in}}}+{Z_{0}}\cos{\theta_{\textrm{tr}}}}},$ (7)
and
$\Gamma_{\textrm{h}}=\frac{{Z\cos{\theta_{\textrm{tr}}}-{Z_{0}}\cos{\theta_{\textrm{in}}}}}{{Z\cos{\theta_{\textrm{tr}}}+{Z_{0}}\cos{\theta_{\textrm{in}}}}},$ (8)
where $\theta_{\textrm{in}}$ is the incident angle, $\theta_{\textrm{tr}}$ is the transmitted wave angle, $Z_{0}=\sqrt{\frac{\mu_{0}}{\varepsilon_{0}}}$ denotes the free space impedance, and $Z$ is the wave impedance of the reflected material, calculated as
$Z=\sqrt{\frac{{{\mu_{0}}}}{{{\varepsilon_{0}}({n^{2}}-{{(\frac{{\alpha\mathrm{c}}}{{4\pi f}})}^{2}}-j\frac{{2n\alpha\mathrm{c}}}{{4\pi f}})}}}.$ (9)
Here, $\mu_{0}$ and $\varepsilon_{0}$ are the free space permeability and permittivity, respectively. $n$ is the complex refractive index of the smooth surface, and $\alpha$ is the absorption coefficient of the incident surface material.
At low frequencies, the surface of the material is considered primarily as a smooth surface. However, as the frequency increases to the THz band, the wavelength is short enough to be very close to the surface roughness of typical materials. Therefore, in the THz band, the material cannot be considered a smooth surface. The modified reflection coefficient is obtained by introducing a roughness factor $\rho$. The Rayleigh roughness factor can be expressed as [145]
$\rho=\exp(-\frac{1}{2}{(4\pi{\sigma_{h}}\frac{{\cos{\theta_{\textrm{in}}}}}{\lambda})^{2}}),$ (10)
where $\sigma_{h}$ denotes the standard deviation height of the material surface roughness. Then, the modified reflection coefficient $\Gamma^{\prime}$ is expressed as
$\Gamma^{\prime}=\rho\Gamma.$ (11)
The modified reflection coefficient model is a function of frequency and angle of incidence, and in reference [79], this phenomenon was verified by experiments. As an example, based on the modified Fresnel model, the frequency-angle two-dimensional reflection coefficient modeling for vertical polarization is proposed. The model of non-metallic materials is written as:
$\begin{split}\Gamma_{\mathrm{v,FA}}=&e^{(-10^{a}f^{2}{\rm{cos}}^{2}\theta_{\textrm{in}})}\\ &\cdot\left(\frac{{\rm{cos}}\theta_{\textrm{in}}-\sqrt{1+\frac{10^{b}}{10^{c}-df^{2}-jf}-{\rm{sin}}^{2}\theta_{\textrm{in}}}}{{\rm{cos}}\theta_{\textrm{in}}+\sqrt{1+\frac{10^{b}}{10^{c}-df^{2}-jf}-{\rm{sin}}^{2}\theta_{\textrm{in}}}}\right),\end{split}$ (12)
where $a$, $b$, $c$, $d$ are the parameters to be fitted, and their means are shown in [79]. The results for a plasterboard are shown in Fig. 5. The analysis revealed a low root-mean-square (RMS) error which is around 0.1 for the materials of glass, tile, board, and plasterboard, indicating a fit between the points of measured data and the modified reflection coefficient model. As observed, the reflection coefficient of the material and its growth rate increased with increasing angle of incidence. Therefore, the model characterizes the reflection coefficient as a function of the frequency and angle of incidence.
Figure 5: Reflection coefficient measurement data and fitting results for plasterboard from 220 to 320 GHz [79].
III-B2 Path loss
THz signals transmitted over channels have distinct characteristics compared to lower frequencies, leading to novel and urgent challenges [43]. Representatively, in wireless propagation, the spreading loss increases quadratically with frequency. Besides, with the THz band operating at exceptionally high frequencies, signals transmitted over a THz link are severely affected by the spreading loss [17]. The channel characteristics of high propagation loss can be characterized by path loss. A novel index is essential for characterizing and modeling the path loss in THz channel. Modeling of path loss in the THz band follows the 5G modeling methods, such as the FI model, ABG model, and CI model shown in (1), (3), and (2), respectively, in Section II.A. Based on the references in Table I, the PLE of the CI model $\alpha_{\mathrm{CI}}$ for indoor and outdoor scenarios is summarized, as shown in Fig. 6. The PLE of the THz band is close to 2 in most cases, especially in LOS scenarios. In addition, the PLEs of the NLOS scenarios are markedly higher than those of the LOS scenario, even up to 4.6. This is because the blocking losses caused by walls, cars, etc., can be as large as tens of dB in NLOS scenarios, which means that severe attenuation can occur.
Figure 6: Representative results of PLE based on measurements of THz channels.
Due to the wide span of the THz band, it will show a stronger frequency dependence. To further investigate the frequency dependence of path loss in THz bands, [17] used the ABG model to fit the measured path loss. The measured path loss and ABG model are plotted in Fig. 7. The frequency dependence factors $\gamma_{\mathrm{ABG}}$ and distance dependence factors $\alpha_{\mathrm{ABG}}$ are 2.1 and 1.93, respectively. This denotes that the frequency dependence is slightly stronger than that in the free space [43].
Figure 7: Multiple-frequency measurement data and path loss modeling results from 220 to 330 GHz [17].
III-B3 Channel sparsity
When THz waves impinge upon “rough” surfaces, they give rise to diffuse scattering phenomena. In comparison to low-frequency specular reflections, diffuse scattering-generated multipath energy is weaker. This leads to a considerable reduction in the quantity and power of NLOS MPCs received at the receiver, resulting in most power concentrated on a few MPCs, namely, sparsity.
The channel sparsity can be expressed in both the angular and delay domains and in the low rank of the channel matrix, which refers to a small number of coefficients containing a large proportion of the energy [146]. It is often claimed that mmWave channels are “sparse,” i.e., have few entries in the delay angle bins (ITU-R M.2412 [42]). The sparse mmWave channel at 28 GHz has been discovered and studied in 2018 [147]. So, due to the weak diffraction ability, the THz channel is more likely to be sparse due to the weak diffraction ability [16, 148]. The THz channel is more likely to be sparse due to the weak diffraction ability [16, 148]. [149] quantified the sparsity of the channel using the Gini index. Channel measurements conducted in sub-6 GHz, mmWave, and THz bands confirmed the sparsity characteristic in the THz channel. Then, an intracluster power allocation model within clusters was proposed to model the sparsity in the delay domain. [72] analyzed the number of clusters of delay, angle and power domains. The cluster numbers were reduced, which reflected the sparsity of the THz channel.
III-C THz Channel Modeling
Existing THz channel modeling approaches can be broadly classified as stochastic, deterministic, and hybrid approaches [150]. THz stochastic channel models are mainly obtained based on measurements and describe the type of environment rather than a specific location. While parameterizing such models from measurements requires significant effort, implementing channel modeling from stochastic models is much less complex than implementing deterministic models [151].
THz deterministic approaches solve (approximate) Maxwell’s system of equations in a given environment and can achieve high accuracy [152], but they require detailed information about the geometric and electromagnetic properties of the environment and have high computational complexity, such as RT method [153]. RT-based THz channel modeling has been extensively utilized in sub-6 GHz and mmWave bands [154]. However, channel modeling using the RT approach for the THz band is different from that in the sub-6 GHz and mmWave bands. These differences result in new opportunities for RT-based channel modeling in THz bands. For the inherent characteristics of RT, RT based on geometric optics can be explained by the high-frequency approximation of Maxwell’s equations. Consequently, the quasioptical characteristics make the RT results in the THz band more reliable and suitable than those in the sub-6 GHz band. Additionally, for THz channel characteristics, the high propagation loss and sparsity characteristics of the THz channel enhance the need for site-specific analysis, which can be well captured by RT simulations.
However, there are still some challenges for RT-based channel modeling in THz bands. Most notably, there is no complete electromagnetic property of the material in THz bands [155]. Additionally, there is a lack of adequate measurements for validating and calibrating RT [148]. For real-time applications, the RT channel model is difficult to implement. Therefore, a data and model dual-driven channel modeling method is presented, which uses channel measurement data to calibrate simulation parameters [26]. The data and model dual-driven channel simulation flow is shown in Fig. 8. The material EM properties in RT simulations are tuned to reach the best agreement in terms of power and delay for the dominant propagation paths. In the beginning, the RT simulation outputs the results corresponding to the initial EM parameters, which are compared with the measurement data. If the results of the power and delay parameters for the dominant paths match (with the objective to minimize the RMS error), the simulation results are output. If the results do not match, the relative permittivity and conductivity of the materials are updated, and the simulation results are output again. This operation is repeated until the best agreement between the simulation and measurement results is achieved.
Figure 8: Data and model dual-driven THz channel modeling flowchart [26].
The performance of the data and model dual-driven THz channel modeling method is demonstrated through a comparison between simulations and channel measurements. Fig. 9 shows the measurement power angle profile in which the five dominant paths are marked by numbers. Furthermore, the delay of the five dominant paths is plotted to perform a comparison. Fig. 9 shows the results of RT simulations for the five dominant paths. By comparing Fig. 9 with Fig. 9, it can be seen that MPCs have good agreement in both the spatial and delay domains between the measurement and RT results. For example, the differences between the measurements and RT simulations for the delay of the five paths are within $\pm$2 ns. The maximum differences in the angle of arrival, angle of departure, delay, and power gain of the dominant paths are $-1.8^{\circ}$, $-1.5^{\circ}$, 0.7 ns and -0.7 dB, respectively. Therefore, the data and model dual-driven channel modeling method can describe well the propagation characteristics, i.e., the delay and spatial dispersion, with reduced simulation complexity in the THz bands [26].
Figure 9: Data and model dual-driven channel modeling method results: a) Power, angle of arrival, and angle of departure measurement results; b) Power, angle of arrival, and angle of departure simulation results [26].
THz hybrid channel models combine deterministic and stochastic channel modeling approaches, providing a balance between accuracy and efficiency, such as in map-based channel models [156] and hybrid ray and graph models [157]. For example, due to the sparsity of the THz channel, the main propagation paths of the THz channel (e.g., LOS path and primary reflection path) are generated by the deterministic model, while the other multipaths are generated by the stochastic model, which ensures the accuracy of the main propagation paths and quickly obtains the remaining multipath information [158]. Therefore, the hybrid channel model is a compromised approach for THz channel modeling.
III-D Summary and Prospects
In summary, extensive THz channel measurement campaigns have been carried out in typical scenarios, such as InH, data centers, and outdoor streets. Existing studies indicate that the THz band exhibits significantly different channel characteristics compared to sub-6 GHz and mmWave bands, including characteristics such as reflection, high path loss, and sparsity. Stochastic, deterministic, and hybrid channel modeling approaches have been employed in THz channel modeling. Among them, the data and model dual-driven THz channel modeling method show promising prospects. Despite the research on THz channels in recent years, many open issues motivate future research efforts in the following areas [25].
III-D1 THz channel measurement platforms with higher performance
Current THz channel measurement platforms are unable to achieve a balance between measurement accuracy, measurement speed, and measurement distance, which constrains the study of THz channel characteristics. In addition, existing virtual MIMO THz channel measurements based on rotating horn antennas are time-consuming. Therefore, it is necessary to build a THz E-MIMO array and phased array to achieve accurate parameter estimation and fast beam sweeping.
III-D2 THz channel characteristics
An increasing number of frequency bands will be deployed and coexist in the future, ranging from sub-6 GHz to 10 THz. In multiband parallel communication systems, it is possible to use the channel characteristics extracted from one frequency band to assist the link establishment of the other band. However, the frequency separation between the bands under consideration is large, and many characteristic channel variations must be considered. Moreover, new characteristics of the THz channel, such as sparsity, near-field characteristics, and nonstationary characteristics, need to be modeled, characterized, and incorporated into the channel model.
III-D3 Frequency-dependent THz channel model
THz bands have a spectral range of nearly 10 THz, while the 5G standard channel model only supports a frequency range of 0.5 to 100 GHz. Therefore, the channel model of THz bands should be frequency-dependent not only for large-scale parameters but also for small-scale parameters to cover such wide frequency bands.
IV E-MIMO Channel Measurement and Modeling
The concept of large antenna arrays was proposed in [159]. In [160], the number of antennas tended to infinity, and favorable propagation was valid under the condition of a practical angle distribution. In the 5G era, 3D MIMO is considered a promising practical technology, as shown in Fig. 3. The 3D channel model is an extension to the 2D channel model by including elevation angles at both the departure and arrival, which makes full use of the elevation domain and increases capacity [67]. It is easier to satisfy favorable propagation conditions with a large angular spread in azimuth and elevation. In addition, the spatial correlation of the 3D channel was theoretically analyzed, which showed that the spatial correlation of the 3D channel is inversely proportional to the vertical space and angular spread between the antenna elements. The capacity of the 3D channel is significantly higher than that of the 2D channel [161]. In Fig. 10, a large capacity gap between $C_{\mathrm{3D}}$ and $C_{\mathrm{2D}}$ can be observed (38.4% at SNR = 25 dB, EAS = 10) and larger EAS values will expand the gap. The derived tight bounds fit well with $C_{\mathrm{3D}}$ for both EAS = 10 and 20. In [162], to further investigate 3D massive MIMO channels, field measurements were made on 32 to 256 antenna units at the transmitter. Based on the channel information extracted from the measurement data, the power angle spectrum, RMS angle spread, and channel capacity were studied. With the number of antennas increasing, the angle distribution is more diffuse and channel capacity increases.
Figure 10: Channel capacity: 3D channel model and 2D channel model [161].
It is well known that massive MIMO technology can significantly increase the wireless transmission data rate, improve the communication quality, and further explore the wireless resources in the space domain. Massive MIMO has more than one hundred antennas on the BS side, forming a large-scale multiantenna array. It makes full use of the spatial degree of freedom, which can effectively improve the transmission capacity and power efficiency of wireless communication systems. In the 6G era, the number of antenna elements in E-MIMO systems is expected to increase compared to massive MIMO and form much larger-scale antenna arrays, such as extremely large aperture array and holographic massive MIMO [6]. In some studies, the number of E-MIMO array elements reached the order of hundreds or thousands [50]. Predictably, with rapid technological innovation and enormous commercial demand, massive MIMO will be further expanded, and the number of array elements will range from hundreds to thousands, forming an E-MIMO array that will change the channel characteristics and models [26, 163]. Meanwhile, some new characteristics will be introduced, including near-field effect and SnS characteristics. Therefore, this will likely completely reshape the future 6G communication field.
IV-A E-MIMO Channel Measurement
The methods for MIMO channel measurement are divided into two categories. One is to use a virtual antenna array (VAA), and the other is to use a real antenna array (RAA). Table II summarizes some channel measurements. Only a few channel measurement activities have employed RAA for measurements [164, 165, 166] in Table II. However, this method requires an expensive investment and must take into consideration the problem of massive MIMO antenna calibration. Therefore, the VAA is more common in measurements, especially for E-MIMO with thousands of antenna array elements, as shown in Table II. For VAA measurements, there are two ways to form a VAA. One method is an antenna array composed of a single custom antenna with a mechanical 3D turntable, which has been adopted in many studies due to its convenience and easy operation [167, 168, 163]. The second method is to use a small antenna array, which is conducive to splicing more array elements in a shorter time, and the measurement quantity is relatively smaller [162], as shown in Fig. 11. In the measurement with a virtual antenna, there are some problems, including the synchronization of the transceiver and receiver and the phase deviation caused by array movement.
In addition to the antenna configuration, there are two main channel measurement platforms, including a frequency domain measurement platform based on a VNA and a time domain measurement platform based on a time-impulse sounder. The measurement system based on VNA has the advantages of high time domain resolution, internal synchronization of the transmitter and receiver, and low complexity. However, this method must connect the transceiver with the VNA, which limits its application in remote outdoor scenarios. To solve this problem, some researchers have proposed prolonging the measurement distance of VNAs by taking advantage of the low loss characteristics of optical fibers [169]. This system is usually used for E-MIMO channel measurement in InH scenarios, including conference halls and auditoriums. For the time domain measurement system, the correlation-based channel measurement system has the advantages of instantaneous broadband measurement, fast measurement, and direct acquisition of time domain results. Although a wired connection between Tx and Rx is not required in this method, synchronization is required between Tx and Rx [67, 170]. The system may be suitable for outdoor long-distance E-MIMO channel measurements. In [171], a larger-scale configuration of two antenna arrays was proposed for monitoring the birth and death of MPCs by comparing the frequency correlation signature of the arrays, thus enabling a reduction in the density of spatial sampling.
Figure 11: A virtual massive MIMO channel measurement [162].
TABLE II: Massive MIMO and E-MIMO measurements and antenna configurations
Number of array elements Frequency (GHz) Antenna Channel characteristics Reference
Tx: 16; Rx: 64 5.8 RAA
Near-field: Power, condition number, etc.
[164]
Tx: 128; Rx: 1 2.6 VAA, RAA
Near-field: Capacity, PAS,
singular value spreads, etc.
[172, 165]
Tx: 4; Rx: 128 5.3 RAA
Near-field: Correlation,
singular value, etc.
[166]
Tx: 1; Rx: 128 5.6
VAA
Near-field: DS
[167]
Tx: 128 or 64; Rx: 1 or 4 26 VAA
Near-field: PDP, DS,
coherence bandwidth, etc.
[168]
Tx: 256; Rx: 16 3.5 VAA
Far-field: PDP, DS, etc.
[170]
Tx: 256; Rx: 16 3.5, 6
VAA
Far-field: PAS, AS,
capacity, etc.
[162, 173]
Tx: 1; Rx: 256 11 VAA
Near-field: PAS, power, cluster lengths,
cluster arrival interval, etc.
[174]
Tx: 21$\times$21; Rx: 3$\times$3 26-30 VAA
Far-field: PDP, DS, AS, etc.
[175]
Tx: 720; Rx: 1 26.5-32.5 VAA Near-field: CIR, power [176]
Tx: 1; Rx: 10$\times$10$\times$10 5.15 VAA
Near-field: Spherical wave coefficient
[177]
Tx: 1; Rx: 1600 15 VAA
Near-field: K,
DS, AS
[163]
Tx: 1; Rx: 1800 26-30 VAA
Near-field: CIR, PADP
[171]
Tx: 1; Rx: 2400 100 VAA
Near-field: Delay, PAS,
path gain, etc.
[26]
RAA: real antenna array, VAA: virtual antenna array
IV-B E-MIMO Channel Characteristics
Table II shows the analysis of channel characteristics in near-field and far-field region, including power, delay spread, power angular spectrum, etc. In addition, these channel characteristics exhibit the SnS along the array.
IV-B1 Near-field effect
With the largest phase discrepancy among all BS and user sider antennas reaching $\pi/8$, the distance between the center of a BS array and the center of a user side array is defined as the Rayleigh distance. In general, the Rayleigh distance of an antenna array is defined by $2D^{2}/\lambda$, where $D$ and $\lambda$ are the sizes of the array and carrier wavelength, respectively [178], also known as the Fraunhofer distance. Since the number of antennas is not very large in the 5G massive MIMO system, the Rayleigh distance is on the level of a few meters, which is negligible. Therefore, the existing 5G communication is mainly developed from far-field communication theory and technology. However, with the significant increase in the number of antennas and carrier frequencies in future 6G systems, the near-field region of E-MIMO will be expanded by several orders of magnitude. For example, a 7.3-meter-long virtual linear array was adopted in [165], and the corresponding Rayleigh distance was over 900 m at 2.6 GHz, which is much larger than the radius of a typical 5G cell. Therefore, near-field MIMO communication will become the basic component of the future 6G mobile network, where an accurate channel model considering near-field propagation is necessary, such as spherical wave model.
In [179], the sizes of the antenna Fresnel and Fraunhofer field regions were systematically derived starting from a general phase factor representation of the scalar diffraction theory. However, with the introduction of novel technologies, the near-field range in RIS systems was determined by the harmonic mean of the BS-RIS distance and RIS to user equipment (RIS-UE) distance [48]. In addition, as the antenna array size becomes larger with a large number of antenna elements, the distance between the receiver and transmitter may be shorter than the Rayleigh distance, and the far-field and plane-wave front assumptions in the traditional channel model no longer apply to E-MIMO. The spherical wavefronts should be taken into consideration [180]. The plane wave can be considered a long-distance approximation of a spherical wave. In the far-field region, the phase of the electromagnetic wave can be approximated by a linear function of the antenna index by Taylor expansion. In [48], with plane wavefronts, far-field beamforming can redirect the beam energy to specific angles at different distances. In the near-field region, the phase of the spherical wave is precisely derived based on physical geometry and is a nonlinear function of the antenna index. The information on the incident angle and distance in each path between the BS and UE is embedded in this nonlinear phase. In addition, in the near-field region, E-MIMO can present non-stationary characteristics on arrays [176, 181], which are described in the following section.
IV-B2 Spatial nonstationarity
Since E-MIMO systems are equipped with large-scale antenna arrays, the different regions of the array see different propagation paths, resulting in some channel characteristics showing nonstationarity on the array, which is called the SnS characteristic. In [172], both linear and cylindrical arrays experienced large-scale fading over the array. In Fig. 12, the SnS characteristic is shown in the E-MIMO channels since the objects in the scenario may no longer serve as complete scatterers for the entire antenna array as its aperture increases. The SnS characteristic of multipaths has been observed in E-MIMO channel measurements [176]. The different channels are shown by the elements marked in green and the reference element in Fig. 12. Therefore, unlike previous MIMO systems, the SnS characteristic is a new challenge in E-MIMO scenarios and must be considered in channel modeling. The different antennas at the BS can observe different clusters at different times, which can be modeled as a birth-death process [182]. However, for the existing statistical channel models, channel nonstationarity is not considered, or the characterization and interpretation of SnS are not sufficient. Therefore, channel nonstationarity should be handled in the E-MIMO channel model for 6G [16].
Figure 12: Spherical propagation with SnS characteristics in the near-field region. Three cases that contribute to SnS propagation, i.e., LOS blockage, incomplete reflection, and diffraction, are illustrated [176].
Figure 13: SnS measurement and simulation: (a) Rx view in an empty-room scenario; (b) Ray trajectories; (c) Power-delay-element measurement result; (d) Power-delay-element simulation result [26].
IV-C E-MIMO Channel Modeling
The assumption of a plane wave and channel spatial stationary is used in the 3GPP 38.901 channel model of 5G, which is mainly aimed at small-scale MIMO systems [43]. However, these assumptions will be challenged in massive MIMO systems, especially E-MIMO systems. The Rayleigh distance of E-MIMO will become larger in a real deployment scenario. In this case, the UE or scatterer is most likely to be in the near field of the BS. Thus, the more general spherical wave model should be considered for channel modeling.
Statistical channel modeling is an important modeling approach for massive MIMO channel modeling, which can be further divided into the correlation-based stochastic model (CBSM) and geometry-based stochastic model (GBSM). Due to the low complexity and limited precision, the CBSM is not sufficient to simulate the above non-stationary phenomena or spherical wave effects. The GBSM is more practical because of its relatively higher accuracy. In [183], second-order approximate channel models of spherical wavefronts (i.e., parabolic wavefronts) for massive MIMO channels were established in both spatial and temporal domains to effectively simulate near-field effects. In [184], the general 3D non-stationary GBSM for E-MIMO communication systems was proposed in 6G. Based on the complex principal component analysis (PCA) method in machine learning, the principle involves maximizing the converted channel power, extracting representative channel characteristics, and using these characteristics to ultimately reconstruct the CIR [185]. The proposed method is robust with more antennas, which provides insights for future E-MIMO channel modeling. In addition, the COST 2100 model is a cluster-level GBSM for the MIMO system, which is the first to use the concept of the visible region (VR) [186] to simulate SnS along the antenna array. The cluster can only be viewed through the antenna element located in the corresponding VR. However, due to the plane wave hypothesis, the modeling precision of SnS characteristics is limited in the COST 2100 model. Based on the VR concept, the GBSM with different frequency bands and scenarios is used to represent SnS and near-field effects for the large-scale array configuration [174]. However, the validity of these models needs to be further verified by more measurements in real scenarios.
To reduce the complexity of E-MIMO channel modeling, [176] proposed a channel modeling framework based on the capture and observation of multipath propagation mechanisms (i.e., LOS, reflection, and diffraction). In Fig. 13, it is shown that spherical propagation with the SnS characteristic is based on measurement and RT simulation. Fig. 13(a) shows the measurement environments. Fig. 13(b) shows the ray trajectory diagrams of the dominant path. The trajectory color changes in the order of red, orange, yellow, and green, indicating that the received power decreases sequentially. Compared with traditional statistical channel modeling, only one additional SnS parameter is added to the proposed framework, which is required for low-complexity implementation. This method has good real-time performance, low complexity, and accuracy. In [176], there are $K$ SnS spherical propagation paths between the Tx array and Rx. The E-MIMO channel at frequency $f$ can be modeled as a superposition of channel frequency responses (CFRs) of the $K$ paths on the array, which is Fourier transform of CIR and can be expressed as
$\begin{split}\mathbf{H}^{\rm sns}(f)&=\mathbf{S}\odot\mathbf{A}(f)\cdot\mathbf{H}(f),\end{split}$ (13)
where $\mathbf{H}^{\rm sns}(f)$ comprises $M$ complex values, i.e., $\mathbf{H}^{\rm sns}(f)\in C^{M\times 1}$, $f\in[f_{L},f_{U}]$ is the frequency within the designed range, and $\odot$ represents the elementwise product operation. $\mathbf{H}(f)\in\mathbb{C}^{K\times 1}$ denotes CFRs at $f$ of the $K$ paths at the reference point (in Fig. 12).
$\begin{split}\mathbf{H}(f)&=[\alpha_{1}e^{-j2\pi f\tau_{1}},...,\alpha_{k}e^{-j2\pi f\tau_{k}},...,\alpha_{K}e^{-j2\pi f\tau_{K}}]^{T},\end{split}$ (14)
where $\alpha_{k}$ and $\tau_{k}$ represent the complex amplitude and propagation delay of the $k$th path, respectively. $(\cdot)^{T}$ denotes the transpose operation. $\mathbf{A}(f)\in\mathbb{C}^{M\times K}$ is the array manifold matrix. The manifold projected on the $m$th antenna element by the $k$th path, i.e., $\mathbf{A}$’s $\left(m,k\right)$th entry $a_{m,k}$, can be represented by the transfer difference of the $m$th element with respect to the reference point:
$a_{m,k}\left(f\right)=\frac{\left\|\bm{d}_{k}\right\|}{\left\|\bm{d}_{m,k}\right\|}e^{-j2\pi f\frac{\left\|\bm{d}_{m,k}\right\|-\left\|\bm{d}_{k}\right\|}{\textbf{c}}},$ (15)
[…]
A novel matrix $\mathbf{S}$ is introduced in the proposed modeling framework for the SnS characteristic. It contains $K$ nonnegative real-valued vectors, i.e.,
$\begin{split}\mathbf{S}&=[\textbf{{s}}_{1},...,\textbf{{s}}_{k},...,\textbf{{s}}_{K}],\end{split}$ (16)
where $\textbf{{s}}_{k}=[\textit{s}_{1,k},...,\textit{s}_{m,k},...,\textit{s}_{M,k}]^{T}$ and $\textit{s}_{m,k}$ is the SnS property of the $k$th path on the $m$th element. In Fig. 13(c) and (d), the power and delay of the MPCs vary with the element, exhibiting the SnS characteristic. In addition, the delay differences between the three dominant paths are 0.1 ns, and the power differences are within 0.3 dB, which verifies the accuracy of the channel modeling method.
Compared with statistical channel modeling, deterministic channel modeling can accurately capture electromagnetic wave propagation characteristics. The RT method can accurately capture channel characteristics along E-MIMO antenna arrays, which is a promising channel modeling method for E-MIMO systems [187, 188]. However, only a few studies have implemented RT to characterize channels with large array configurations. The reason is that the arrays on the Tx and Rx sides are usually obtained by strong simulation of each Tx and Rx antenna pair in RT. The computational complexity is particularly challenging for E-MIMO systems because it increases rapidly with the number of array elements. Recently, the map-based model has been proposed to implement the RT method, which can effectively reduce the computational complexity by simplifying the geometry and electromagnetic description of the environment and restricting the sequence of interactions between rays and objects [189]. The model may have poor accuracy in predicting the channel at specific locations, but it is fully capable of simulating the non-stationarity and transition of large arrays. However, the strategy has difficulty capturing SnS between array elements (i.e., ray birth-death process). In addition, the hybrid modeling approach that combines statistical and deterministic modeling is expected to reduce the complexity of the model because only the dominant path has RT characteristics, but the key issue is the simplicity and accuracy of channel modeling.
IV-D Summary and Prospects
The E-MIMO channel measurement mainly adopts the frequency domain or time domain channel measurement platforms based on the VAA. With the increase in the 6G spectrum, near-field effects and spatially non-stationary characteristics are more likely to appear. To address these characteristics, scientists have carried out statistical and deterministic modeling, and some achievements have been made. However, there are still many unsolved problems and challenges. This section summarizes the channel research method, new characteristics, and modeling method and discusses several future research directions.
IV-D1 Channel measurements for more scenarios and bands
There are many dynamic application scenarios in 6G, but the current VAA measurement method has difficulty meeting the measurement requirements of the E-MIMO system in a dynamic environment, and the phase sensitivity of large array antennas should be considered. Besides, 6G has more frequency bands, including mid-band, mmWave, and THz and so on. Channel measurements with larger antenna arrays are needed at these frequency bands.
IV-D2 Channel characteristics combined with new technology
SnS is an important channel characteristic of E-MIMO. Improving the description of the SnS characteristic on arrays is also an important topic in 6G channel research. In addition, the impact of novel technologies in near-field communication (such as RIS-assisted near-field communication, THz communication) has not been fully studied. For some performance metrics, such as capacity, the classical Rayleigh distance may not capture the performance loss of these metrics well.
IV-D3 Channel modeling based on ITU/3GPP framework
In real systems, both far-field and near-field signals are usually present in the communication environment. Compared with the existing GBSM model based on plane waves, the spherical wave model can be introduced to solve the near-field effect, but the model is too complicated. The deterministic modeling approach can accurately capture massive MIMO channel characteristics, but the computational complexity may be high. To balance the relationship between complexity and accuracy or to choose channel modeling methods according to the needs of the system, corresponding solutions should be provided in future studies.
V ISAC Channel Measurement and Modeling
Figure 14: Illustration of the ISAC channel model and new characteristics. The dashed lines indicate that the paths are interrupted[54].
At the 44th meeting of ITU-R Working Party 5D (WP 5D), ISAC officially became one of the six typical application scenarios of 6G. It introduces sensing into communication systems, utilizing the transmission, reflection, and scattering of radio waves to understand the physical world. In communication systems, only the communication channel is considered. For better illustration, Fig. 14 shows the channel model of the ISAC system to understand the differences between communication and ISAC channels. This channel contains a typical ISAC scenario with one BS and one mobile terminal. In the subsequent descriptions, Tx is used to describe BS. Rx_S and Rx_C are used to describe the receiving devices of sensing and communication, respectively. When Rx_S and Tx are in the same position, the sensing signals are scattered by scatterers and returned to Rx at the same angle, creating an echo channel. This mode is defined as mono-static sensing. However, when Rx_S and Tx are not in the same position, it is referred to as a bi-static sensing mode. The model takes mono-static sensing as an example: Tx transmits sensing and communication signals and receives sensing signals, while the mobile terminal is designed to receive communication signals.
V-A ISAC Channel Measurement Progress
Channel measurement activities are a key step in conducting ISAC channel research, through which new channel characteristics can be discovered and regular conclusions can be summarized. Many studies have measured ISAC channels that focus on various perspectives, such as frequency, scenario, and sensing modes. At the same time, analysis has been conducted on the characteristics of the ISAC channel, such as path loss, angular spread, delay spread, and RCS. Because different sensing modes lead to different measurement methods, the measurement activities of ISAC channels are summarized in terms of mono-static and bi-static sensing.
To simulate mono-static sensing, the Tx and Rx-S are placed in the same position during channel measurements, using physical isolation to eliminate the mutual influence of the Tx and Rx-S antennas. For instance, a research team uses this approach to scan the surrounding environment while steering the beam pattern toward different directions and observing the target reflections [190]. Moreover, ISAC measurements are conducted to investigate and model the path loss in the sensing channel under the mono-static sensing mode[191]. A measurement activity measures the sensing channel and finds that the received power is affected by the sensing target, and the Tx to Rx_S and Rx_S to Tx channels will produce self-interference, which should be portrayed in the ISAC channel model [192]. The range-power profile, range-Doppler spectrum, and range-angle spectrum are obtained and analyzed in mono-static sensing measurements [193, 194].
Another mode of sensing is bi-static sensing, where Tx and BS_S are placed in different positions to sense the environment or target. In [195], the cluster power attenuation factor, which is a new parameter of the ISAC channels, is measured in the bi-static sensing mode. Small-scale fading is one important characteristic of the channel, which is measured and modeled [196]. The Doppler characteristics under the bi-static sensing mode are studied and modeled based on measurement data[197]. To reconstruct the indoor environment, some researchers combine the angle of departure and angle of arrival information of the sensing channel[198]. Based on the above measurements, researchers begin to conduct preliminary research on sensing channel characteristics under different sensing modes. However, the current measurements primarily focus on recognizing traditional features, such as path loss, PDP, angular spread, and delay spread. There is a lack of new features that characterize the correlation between sensing and communication channels in ISAC channels.
A detailed summary of the measurement campaigns is shown in Table III, which outlines the sensing modes, scenarios, and channel characteristics. The table III provides several key findings. First, there are many studies focused on the measurement of ISAC channels in mm-waves. Because the onboard radar mainly focuses on millimeter waves, the wavelength is short and the bandwidth is large, which is conducive to high-precision perception and high-speed communication. Second, for the measurement methods, the frequency domain VNA and correlation-based time-domain channel platforms are the two primary methods to measure the ISAC channel [52]. The pros and cons of the two methods have been well compared and explained in Section III, and this section does not elaborate further. Third, measurement scenarios are mainly divided into indoor and outdoor scenarios, but detailed scenarios are related to the application of ISAC [51]. For example, for gesture sensing, it is necessary to construct finger models and conduct research on specialized scenes, and for vehicle-pedestrian detection, measurements need to be made on the street.
TABLE III: ISAC channel measurement campaigns
Sensing mode
Measurement
platforms
Frequency
/BW (GHz)
Scenarios Channel characteristics Reference
Mono-static 0.1-3/2.9 Foliage environment
PL, PDP, SF,
inter-cluster parameters
[191]
Correlation-
based
28/1 Indoor office PL, Doppler, RCS [190]
70.6/2 Indoor empty room Power, delay, Doppler, RCS [192]
73/2 V2V Doppler, RCS, power [193]
VNA 57.5-63.5/6 Indoor lab PL, RCS, reflection [194]
Bi-static Correlation- based 24/0.25 Laboratory
Doppler, PL, RCS,
SnS
[197]
60/0.1 Corridor Delay, AS, PL, DS [198]
VNA 2-8/0.3 Anechoic chamber
Inter-cluster parameters,
PL, AS, DS
[196]
8-10
Meeting room,
classroom
SnS, PDP, near-field,
intra-cluster parameters
[195]
V-B ISAC New Channel Characteristics
The ISAC channel model needs to consider both sensing and communication services, and there are some new features emerging in terms of channel characteristics, posing a big challenge to channel modeling. For example, sensing is to obtain the characteristics of the target, so the channel model needs to include the target, which is not present in communication channels. There may be some multipath overlap between the sensing and communication channel in the environment, and it is necessary to study their sharing characteristics. This section lists the new channel characteristics of the ISAC channel.
V-B1 Target scattering
Because the purpose of sensing is to obtain information about a target, target scattering parameters play an important role in sensing channel modeling. In the existing research, for modeling the target scattering, the RCS is the main parameter to describe the target in the sensing channel, which presents the interception and scattering ability of the target to the signal power. When the signal is incident at different angles, the scattering ability of the target varies, affecting the power of the sensing signal. In general, the RCS modeling is used to measurements or deterministic methods, such as approximate solutions to Maxwell’s equations and measurements in an anechoic chamber. Researchers [199] propose a modeling method for sensing targets, and the method is verified by actual measurements. Most studies on RCS assume that the electromagnetic wave that reaches the target is a plane wave in the far-field conditions. However, in ISAC applications, the LOS path sensing distance is typically several hundred or even tens of meters, resulting in most of the targets being in the near field of the antenna. A comprehensive RCS unity model must be developed, encompassing both near- and far-field properties, to apply this to diverse targets within ISAC systems.
V-B2 Doppler coupling
To achieve high-precision sensing, the ISAC system requires support from high carrier frequencies. However, the Doppler effect, which is proportional to the carrier frequency, becomes even more stringent in ISAC systems, particularly in high-mobility scenarios. Some studies take into account the differences in Doppler parameters of ISAC channels [198]. If ISAC technology is implemented in a widely used orthogonal frequency division multiplexing (OFDM) system, the sensing channel with high Doppler propagation may disrupt subcarrier orthogonality and cause intercarrier interference, while the accuracy of the speed estimate is reduced.
V-B3 Environmental interference
Figure 15: Target response measurements and results: (a) Measurement scenario layout; (b) Power delay profile (PDP) of the result [200].
Figure 16: Measurements and results of shared scatterers: (a) Sensing and communication measurement environment in an indoor hall; (b) Sensing PADP; (c) Communication PADP[54].
Sensing signals are influenced by scatterers besides the target during propagation, thus affecting the recognition of the target. Therefore, in the ISAC channel, it is necessary to consider the interference caused by scatterers, which is called environment interference. The interference of the environment is explored by conducting measurements with and without metal interference on an outdoor road [200]. The measurement scenario layout is shown in Fig. 15(a). The red dots in Fig. 15(a) represent the measurement positions of Scheme 1, where the environment interference on the sensing channel is studied at different distances. The CIRs of each distance are stitched to obtain the measured power delay profile (PDP), determining the ability of the channel to sense the target from the environment. The purple dots and stars represent the stationary sensed target and the moving metal interference of Scheme 2, respectively, to investigate the effect of the interference on the sensing channel at different distances. When the metal approaches the target, other multipath paths are generated near the target path, as shown by the lines labeled 7 and 8 in Fig. 15(b), thus affecting the target recognition. To quantify the variant of the environment around the sensed target and the proportion of the LOS path in the multipath, the power ratio (PR) is defined, which is the ratio of the sensed target power and the sum of the effective environmental power. The results indicates that when the environment is complex, the PR value decreases, indicating that the multipath of the target is difficult to identify. In addition, the distance factor has an impact on the PR value, and the PR value decreases as the distance increases. However, considering interference into the ISAC channel model still needs to combine with additional channel measurement analysis and simulation verification.
$h_{c}(\theta_{c,R},\theta_{c,T},\tau_{c})=$ $\underbrace{\sum\limits_{n_{0}=1}^{N_{0}}\sum\limits_{m_{0}=1}^{M_{0}}a_{c,n_{0},m_{0}}\delta(\theta_{c,R}-\theta_{c,n_{0},m_{0},R})\delta(\theta_{c,T}-\theta_{c,n_{0},m_{0},T})\delta(\tau_{c}-\tau_{c,n_{0},m_{0}})}_{\text{Shared\ Communication\ Sub-Clusters}}$
$+\sum\limits_{n_{1}=1}^{N_{1}}\sum\limits_{m_{1}=1}^{M_{1}}a_{c,n_{1},m_{1}}\delta(\theta_{c,R}-\theta_{c,n_{1},m_{1},R})\delta(\theta_{c,T}-\theta_{c,n_{1},m_{1},T})\delta(\tau_{c}-\tau_{c,n_{1},m_{1}}),$ (17a)
$h_{s}(\theta_{s,R},\theta_{s,T},\tau_{s})=$ $\underbrace{\sum\limits_{n_{0}=1}^{N_{0}}\sum\limits_{m_{0}=1}^{M_{0}}a_{s,n_{0},m_{0}}\sigma_{n_{0},m_{0}}\delta(\theta_{s,R}-\theta_{s,n_{0},m_{0},R})\delta(\theta_{s,T}-\theta_{s,n_{0},m_{0},T})\delta(\tau_{s}-\tau_{s,n_{0},m_{0}})}_{\text{Shared\ Sensing\ Sub-Clusters}}$
$+\sum\limits_{n_{2}=1}^{N_{2}}\sum\limits_{m_{2}=1}^{M_{2}}a_{s,n_{2},m_{2}}\sigma_{n_{2},m_{2}}\delta(\theta_{s,R}-\theta_{s,n_{2},m_{2},R})\delta(\theta_{s,T}-\theta_{s,n_{2},m_{2},T})\delta(\tau_{s}-\tau_{s,n_{2},m_{2}}).$ (17b)
V-B4 Sharing feature
In the environment, a segment of scatterers introduces perturbations that can exert a notable influence on the propagation of communication signals, thus potentially impairing transmission quality. Furthermore, these scatterers might concurrently impact the accuracy of target detection and recognition processes. The overlapping scatterers contributing to both communication and sensing channels are defined as shared scatterers, as shown in Fig. 14. These shared scatterers are validated and analyzed through realistic ISAC channel measurements conducted in [54]. Fig. 16(a) illustrates the layout of the ISAC channel measurement scenario, with primary scatterers identified by red sequence numbers. Based on the CIRs obtained through measurements, the power angular delay profiles (PADPs) of sensing and communication channels are calculated, as shown in Fig. 16(b) (c). As shown in Fig. 16, the sensing multipaths corresponding to the scatterers exhibit a realistic environment, while the communication multipaths indicate the dominant scatterers that affect communication signal propagation. A comparison of Fig. 16(b) with Fig. 16(c) shows that some scatterers are shared between the sensing and communication channel, i.e., the south wall, north wall, and east wall. Moreover, the shared clusters generated by these shared scatterers reveal realistic sharing of communication and sensing channels, which need to be adequately considered in channel modeling for ISAC systems.
V-C ISAC Channel Modeling
To model the channel for the ISAC systems, it is necessary to extend the channel modeling method of the above 5G system. First, the ISAC channel model needs to include RCS, which affects the path loss and can be introduced into the large parameter part of the model. Second, the Doppler parameter needs to be extended based on the communication model. Third, to determine the environmental interference, some relevant parameters are introduced, such as PR, into the ISAC channel model for characterization during small-scale modeling. For the sharing feature of the ISAC channel, a shared cluster-based stochastic channel model is proposed [54]. This model involves the superposition of shared and non-shared clusters and includes RCS features. The sharing degree (SD) metric is introduced to measure the PR of the shared clusters in the channels. Some literature analyze the sensing and communication channels based on similar AoD [201]. In addition, a calculation method of ISAC channel based on sensing and communication clusters is proposed [202].
As illustrated in Fig. 14, the clusters are contributed by the shared scatterers in both the communication and sensing channels are defined as the shared clusters of ISAC channels. The specific formulas are given in (17) and (17). In this model, $h_{c}(\theta_{c,R},\theta_{c,T},\tau_{c})$ and $h_{s}(\theta_{s,R},\theta_{s,T},\tau_{s})$ denote the CIRs of communication and sensing, respectively. The subscripts $c$ and $s$ denote the communication and sensing propagation, respectively. $a$ and $\tau$ are employed to describe the complex amplitude and time delay of communication and sensing MPCs, respectively. $\theta$ denotes the azimuth transceiver angle of the MPCs, where $\theta_{T}$ represents the angle of departure and $\theta_{R}$ represents the angle of arrival. A detailed explanation of this formula can be found in [54].
In the proposed ISAC channel model, the sensing and communication channels are generated independently, but they are partially coupled with each other, which provides insights for the subsequent ISAC channel model based on the 5G standard channel model extension. Moreover, due to the changes in channel topology, parameters, such as delay spread and angular spread in the ISAC channel model, need to be supplemented through channel measurement. Moreover, the aforementioned ISAC channel characteristics are taken into account in channel modeling, which is not currently addressed in communication channel models.
V-D Summary and Prospects
ISAC channel measurement primarily focuses on the frequency or time domain, and measurement methods are classified into mono-static and bi-static sensing. With the integration of sensing and communication, RCS, Doppler, environmental effects and shared cluster characteristics are important in an ISAC channel. To capture these characteristics of ISAC channels, a preliminary statistical channel model is constructed [54], and some modeling views are proposed. However, many aspects require further research. This subsection summarizes the current progress of ISAC channel measurement, new characteristics, and modeling and describes several future research directions for ISAC development.
V-D1 New measurement scheme
Channel measurement is an effective method for extracting the characteristics of ISAC channels. Studies primarily analyze the characteristics of path loss, angular spread, delay spread, and RCS for ISAC channel measurements. ISAC channel measurement requires to capture of sensing and communication channels accurately in large bandwidth and high-frequency bands, which presents challenges for generating and receiving two channels simultaneously, storing large amounts of data and calibrating the system. Additionally, the shared integration mode is a big challenge for ISAC channel data acquisition.
V-D2 New characteristics integration
The study of ISAC channel characteristics is based on practical measurement results. Studies on ISAC channel characteristics are primarily focused on traditional communication channel characteristics, such as delay and angle spread, which are inadequate for a comprehensive understanding of the ISAC channel. Therefore, further research is required to explore new characteristics, including shared scatterers, target modeling, clutter modeling, correlation, and other aspects. Moreover, it is necessary to incorporate these features into the ISAC channel model.
V-D3 Channel modeling extension based on 3GPP framework
In terms of ISAC channel modeling, the communication channel model is well developed. Therefore, the extension based on the communication channel model is a mainstream method. The ISAC channel model is supported by supplementing the traditional modeling process and parameters. However, environment reconstruction-based channel modeling [203], map-based hybrid channel modeling [204], and standard extension-based channel modeling are potential research directions in ISAC channel modeling. Moreover, the combination of ISAC channel modeling with other emerging technologies, such as RIS and non-orthogonal multiple access, is a potential direction for future research.
VI RIS Channel Measurement and Modeling
Different from traditional communication technologies, an RIS realizes active control of a wireless signal propagation environment by passively reflecting electromagnetic waves. Fig. 17 shows a typical RIS-assisted communication scenario. In this scenario, the LOS path is obstructed, and an RIS is employed to enhance the received signal. Obviously, traditional channel modeling methods are not suitable for this kind of channel, which requires new research on the RIS channel model. Notably, in Fig. 17, the RIS channel was characterized by a Tx-RIS-Rx cascaded channel, which was composed of a Tx-RIS subchannel, RIS-Rx subchannel, and RIS response. In this cascaded channel, the subchannels required independent modeling and were combined in the RIS response. The subsequent RIS channel modeling described in this paper is based on this method.
Figure 17: Typical RIS-assisted communication scenario [205].
VI-A RIS Channel Measurements
Currently, a few researches on RIS channels based on channel measurements have been carried out. Table IV summarizes some RIS channel measurement campaigns. The current focus of RIS measurement research is on the large-scale characteristics. Most measurement campaigns are aiming at obtaining received signal power to analyze the power gain caused by RIS. In terms of measurement frequency, measurement campaigns are mostly carried out in sub-6 GHz and millimeter band. These measurements are typically conducted in microwave anechoic chambers as well as practical scenarios, with a primary emphasis on the large-scale characteristics of RIS channel. In [206], authors carry out a path loss measurement in a microwave anechoic chamber with three different RISs in different scenarios. This work investigates the free-space path loss in the RIS-assisted broadcasting and beamforming scenarios and compares it with the theoretical model proposed by the author. In addition, in [207], a measurement campaign is carried out in a microwave anechoic chamber to obtain the radiation pattern of RIS.
Some of the other measurement campaigns are carried out in practical scenarios. A near-field RIS channel measurement is carried out in indoor and outdoor environments to research the path loss characteristics of RIS channel [208]. Furthermore, to facilitate channel modeling in system simulations, research [209] proposes a reflective RIS physical model and conducts indoor RIS measurement. This work aims to abstract the RIS physical model for application in practical channel modeling scenarios. In [210], a RIS-assisted channel measurement platform that can be applied to various complex scenarios based on the time domain correlation principle is built. Using this platform, the authors carry out a channel measurement in a typical indoor corridor scenario at 26 GHz, as shown in Fig. 18. They research the path loss characteristic of RIS-assisted channels in corridor scenarios based on the measurement data.
Figure 18: RIS channel measurement in a corridor scenario [210].
TABLE IV: RIS measurement campaigns
Continuous
/Discrete phase
Frequency (GHz) Scenario Channel characteristics Reference
Discrete 5.8 Indoor, outdoor Power gain [208]
Discrete 5.8
Microwave anechoic chamber,
indoor, outdoor
Power gain,
radiation pattern
[207]
Discrete 26 Indoor PL [210]
Discrete 26.9 Indoor Radiation pattern [209]
Continuous 4.25, 10.5 Microwave anechoic chamber Power gain [211]
Continuous 27, 33 Microwave anechoic chamber
Power gain,
radiation pattern
[206]
VI-B RIS Channel New Characteristics
In the RIS channel, there are three types of characteristics that are worth paying attention to, namely the reflection coefficient characteristics of RIS element, RIS radiation pattern characteristics, and RIS cascade loss characteristics. Among them, RIS radiation pattern characteristics consists of RIS overall radiation pattern characteristics and RIS element radiation pattern characteristics. The reasons are as follows.
Firstly, the RIS is one of the most important nodes in the RIS channel, and its response are an important component of the RIS channel. In order to calculate the RIS response, it is necessary to calculate the reflection coefficient of each RIS element in advance. Thus, the RIS element reflection coefficient characteristics are very important. Secondly, the final form of RIS response is RIS overall radiation pattern or RIS elemenmt radiation pattern, therefore, their characteristics can not be ignored either. Fig. 19 shows a feasible process of calculating the RIS radiation pattern, corresponding to (18)-(23). Finally, since the RIS cascaded channel is obtained by cascading Tx-RIS sub-channel, RIS response, and RIS-Rx sub-channel, its loss characteristics are different from traditional communication channel. Thus, RIS cascade loss characteristics are worth studying.
Figure 19: The process of calculating RIS radiation pattern.
VI-B1 Reflection coefficient characteristics
The reflection coefficient characteristics of the RIS element are derived via RIS element hardware models. In the initial research on RIS-assisted communication, the hardware model of the RIS element was typically the ideal regulator model [212]. This implies that the RIS element has the same reflection coefficient regardless of the angle and polarization of the incident signal. However, by comparing it with the Fresnel reflection coefficient, it can be found that the incident angle and polarization dependence of the RIS element reflection coefficient may be ignored. The Fresnel reflection coefficient is widely used to characterize the reflective properties of 3D materials. Generally, the Fresnel reflection coefficient is related to the polarization and angle of the incident signal, as demonstrated by (9) and (10). In fact, for the 2D RIS element, its reflection coefficients also exhibit dependence on the incident angle and polarization.
To provide a more accurate characterization of the reflection coefficient characteristics of the RIS element, researchers have proposed electromagnetic hardware models, including the transmission line model[213], equivalent circuit model[214, 212], generalized sheet transmission conditions (GSTCs) model [215], and load impedance model [216].
In the transmission line model, the tangential components of the electric and magnetic fields are equivalent to the voltage and current; then, the relationship between the reflection coefficient and incident angle and polarization direction can be deduced according to the Kirchhoff formula[213]. The equivalent circuit model is similar to the transmission line model, where the RIS element is equivalent to a circuit. Since the impedance in the circuit is related to the incident angle, the reflection coefficient is also naturally related to the incident angle [214]. The GSTCs model, commonly used in the metasurface field, employs the surface susceptibilities tensor as a characteristic parameter of the RIS element. By combining the GSTCs condition, the reflection coefficient corresponding to the incident signal with any incident angle and polarization direction can be calculated [215]. The load impedance model considers the RIS element as a thin surface with a certain electric and magnetic impedance. Its principle is shown in Fig. 20; that is, upon excitation of the incident signal, the induced electric and magnetic currents are generated on the RIS element, acting as new radiation sources that radiate electromagnetic waves outward and making the RIS element exhibit a certain reflection coefficient. Since the induced electric and magnetic currents are jointly determined by the electric impedance, magnetic impedance, and the incident signal, the reflection coefficient is affected by the incident angle and polarization direction of the incident signal[216].
In the load impedance model, the reflection coefficient for the vertically polarized signal is written as
$\Gamma_{\textrm{v}}=\frac{-Z_{0}}{2Z_{e}\cos\theta_{\textrm{in}}+Z_{0}}+\frac{Z_{m}\cos\theta_{\textrm{in}}}{Z_{m}\cos\theta_{\textrm{in}}+2Z_{0}}.$ (18)
Correspondingly, if the incident signal is parallel polarized, the reflection coefficient is written as
$\Gamma_{\textrm{h}}=\frac{Z_{0}\cos\theta_{\textrm{in}}}{2Z_{e}+Z_{0}\cos\theta_{\textrm{in}}}-\frac{Z_{m}}{Z_{m}+2Z_{0}\cos\theta_{\textrm{in}}},$ (19)
where $Z_{e}$ and $Z_{m}$ are the electrical impedance and magnetic impedance, respectively.
Figure 20: Load impedance model of RIS [216].
VI-B2 RIS radiation pattern characteristics
Firstly, introduce RIS element radiation pattern characteristics. Currently, there are two commonly used methods for modeling the RIS element radiation pattern, namely, mathematical function fitting and electromagnetic scattering theory calculations. In [206] and [217], the RIS element radiation pattern was fitted by the $\textrm{cos}^{q}\theta_{\textrm{out}}$ function. In [216] and [218], the physical optics method (PO) was used to calculate the radiation pattern of the RIS element. In [208] and [219], the radiation pattern of the RIS element was modeled based on calculating its RCS. In [205], the electromagnetic equivalence principle was employed to calculate the RIS element radiation pattern.
Taking the model presented in [205] as an example, we introduce the radiation pattern modeling process of the RIS element below. Assuming that a signal with both vertical and parallel polarization components is incident on an RIS element at a horizontal angle $\phi_{\textrm{in}}$ and a vertical angle $\theta_{\textrm{in}}$, its electric field is $\mathbf{E}_{i}$, and its corresponding magnetic field is $\mathbf{H}_{i}$. According to the equivalence theorem, the equivalent current $\mathbf{J}_{s}$ and equivalent magnetic current $\mathbf{M}_{s}$ on the RIS element surface are calculated as
$\mathbf{J}_{s}=$ $\hat{\bm{n}}\times(\mathbf{H}_{i}+\mathbf{H}_{r}),$ (20)
$\mathbf{M}_{s}=$ $-\hat{\bm{n}}\times(\mathbf{E}_{i}+\mathbf{E}_{r}),$
where $\hat{\bm{n}}$ represents the normal vector of the RIS element, $\mathbf{E}_{r}$ and $\mathbf{H}_{r}$ represent the reflected electric and magnetic fields, which are obtained from the reflection coefficient of the RIS element. We select an observation point $o$ in the direction of $\phi_{\textrm{out}},\theta_{\textrm{out}}$, and the scattering field $\mathbf{E}_{s}$ at point $o$ can be calculated by calculating the radiation of electric and magnetic currents. By removing the effect of the distance of the observing point, the radiation pattern of the RIS element can be obtained as
$F_{\textrm{element}}^{p_{1}p_{2}}(\phi_{\textrm{in}},\theta_{\textrm{in}},\phi_{\textrm{out}},\theta_{\textrm{out}})=\frac{4\pi}{\lambda}\frac{r}{e^{-j\frac{2\pi}{\lambda}r}}\frac{(\mathbf{E}_{s})_{p_{2}}}{(\mathbf{E}_{i})_{p1}},$ (21)
where $r$ represents the distance between the center of the RIS element and point $o$. $\frac{4\pi}{\lambda}$ is caused by the conversion between the electric field and the amplitude of the path. $(\mathbf{E}_{s})_{p_{2}}$ represents the $p_{2}$ polarization component of $\mathbf{E}_{s}$. ${p_{1}p_{2}}$ represents the combination of different polarizations of the incident and reflected waves. For example, $F_{\textrm{element}}^{vh}$ refers to the horizontal polarization component of the reflected wave excited by the vertical polarization component of the incident wave at the RIS element.
Then, introduce RIS overall radiation pattern characteristics. In the case where transmitter and receiver are located in the far field of RIS, there exists an RIS overall radiation pattern. The RIS overall radiation pattern is obtained by coherent superposition of the radiation pattern of each RIS element and is calculated as
$F_{\textrm{ris}}^{p_{1}p_{2}}(\phi_{\textrm{in}},\theta_{\textrm{in}},\phi_{\textrm{out}},\theta_{\textrm{out}})=$ (22)
$\sum_{x,y}^{X,Y}F_{\textrm{element}}^{p_{1}p_{2}}(\phi_{\textrm{in}},\theta_{\textrm{in}},\phi_{\textrm{out}},\theta_{\textrm{out}})e^{\frac{2\pi}{\lambda}(\mathbf{r}_{\textrm{in}}\cdot\mathbf{d}_{x,y})}e^{\frac{2\pi}{\lambda}(\mathbf{r}_{\textrm{out}}\cdot\mathbf{d}_{x,y})},$
where $X,Y$ is the number of rows and columns of RIS, and $\mathbf{r}_{\textrm{in}}$ and $\mathbf{r}_{\textrm{out}}$ denote the direction vector of the incident and outgoing waves, respectively. $\mathbf{d}_{x,y}$ is the position vector of the $(x,y)^{\textrm{th}}$ RIS element in the panel. If we take the center point of the RIS as the reference point, $\mathbf{d}_{x,y}$ can be expressed as
$\mathbf{d}_{x,y}=\left[\begin{array}[]{c}(x-\frac{1+X}{2})\\ (y-\frac{1+Y}{2})\\ 0\end{array}\right]d_{\textrm{element}},$ (23)
where $d_{\textrm{element}}$ is the interval between RIS elements.
VI-B3 RIS cascade loss characteristics
Multiple studies have been conducted on the RIS cascade loss of the RIS channel based on theoretically derived and measurements. They provide rich RIS cascade loss modeling methods and combined reveal the RIS cascade loss characteristics of the RIS channel.
During a period, the relationship between the RIS cascade loss of the RIS channel and the distances $d_{1}$ and $d_{2}$ is very relevant, where $d_{1}$ is the distance between the transmitter and RIS, and $d_{2}$ is the distance between the RIS and receiver. A theoretical RIS cascade loss model in free space under the far-field condition was proposed in [218]. The theoretical analysis demonstrated that the RIS cascade loss is proportional to $(d_{1}d_{2})^{2}$. Similar conclusions were also obtained in measurements of a corridor scene. Authors [210] carried out a channel measurement in a typical indoor corridor scenario and studied the RIS cascade loss characteristic of the RIS channel in corridor scenarios. The results indicate that a modified version of the multiplicative distance model can effectively fit the measurement data.
Furthermore, some different conclusions can be made. In [206], a general RIS cascade loss model and some special models in free space were proposed. These special models apply to RIS-assisted far-field beamforming, RIS-assisted near-field focusing, and RIS-assisted near-field broadcasting cases. The RIS-assisted far-field beamforming and RIS-assisted near-field broadcasting models were verified by several measurements conducted in a microwave chamber at 27 and 33 GHz. Different from above, the authors found that the RIS cascade loss is proportional to $(d_{1}+d_{2})^{2}$ in the RIS-assisted near-field broadcasting case. This indicates that the RIS cascade loss characteristics of the RIS channel are related to the encoding method of the RIS and depend on whether the terminal is in the far field or near field.
VI-C RIS Channel Modeling
Considering the prevalent use of reflective RIS, it is often placed between the transmitter and receiver, resulting in a cascaded channel known as the Tx-RIS-Rx channel. The main objective of this subsection is to construct an effective cascaded RIS channel model to accurately capture the response characteristics of the RIS and channel properties, while balancing the model complexity and accuracy.
$h_{u,s}^{\textrm{ris}}(t,\tau)=$ $\sum_{n_{1},m_{1}}^{N_{1},M_{1}}\sum_{n_{2},m_{2}}^{N_{2},M_{2}}\sqrt{P_{n_{1},m_{1}}P_{n_{2},m_{2}}}\left[\begin{array}[]{c}F_{\textrm{Rx},u,\theta}\left(\theta_{n_{2},m_{2}}^{\textrm{Rx}},\phi_{n_{2},m_{2}}^{\textrm{Rx}}\right)\\ F_{\textrm{Rx},u,\phi}\left(\theta_{n_{2},m_{2}}^{\textrm{Rx}},\phi_{n_{2},m_{2}}^{\textrm{Rx}}\right)\end{array}\right]^{\mathrm{T}}\cdot\left[\begin{array}[]{cc}\exp\left(j\Phi_{n_{2},m_{2}}^{\theta\theta}\right)&\sqrt{\kappa_{n_{2},m_{2}}^{-1}}\exp\left(j\Phi_{n_{2},m_{2}}^{\theta\phi}\right)\\ \sqrt{\kappa_{n_{2},m_{2}}^{-1}}\exp\left(j\Phi_{n_{2},m_{2}}^{\phi\theta}\right)&\exp\left(j\Phi_{n_{2},m_{2}}^{\phi\phi}\right)\end{array}\right]$ (24)
$\cdot\left[\begin{array}[]{cc}F_{\textrm{ris}}^{\theta\theta}(\phi_{n_{1},m_{1}}^{\textrm{in}},\theta_{n_{1},m_{1}}^{\textrm{in}},\phi_{n_{2},m_{2}}^{\textrm{out}},\theta_{n_{2},m_{2}}^{\textrm{out}})&F_{\textrm{ris}}^{\theta\phi}(\phi_{n_{1},m_{1}}^{\textrm{in}},\theta_{n_{1},m_{1}}^{\textrm{in}},\phi_{n_{2},m_{2}}^{\textrm{out}},\theta_{n_{2},m_{2}}^{\textrm{out}})\\ F_{\textrm{ris}}^{\phi\theta}(\phi_{n_{1},m_{1}}^{\textrm{in}},\theta_{n_{1},m_{1}}^{\textrm{in}},\phi_{n_{2},m_{2}}^{\textrm{out}},\theta_{n_{2},m_{2}}^{\textrm{out}})&F_{\textrm{ris}}^{\phi\phi}(\phi_{n_{1},m_{1}}^{\textrm{in}},\theta_{n_{1},m_{1}}^{\textrm{in}},\phi_{n_{2},m_{2}}^{\textrm{out}},\theta_{n_{2},m_{2}}^{\textrm{out}})\end{array}\right]$
$\cdot\left[\begin{array}[]{cc}\exp\left(j\Phi_{n_{1},m_{1}}^{\theta\theta}\right)&\sqrt{\kappa_{n_{1},m_{1}}^{-1}}\exp\left(j\Phi_{n_{1},m_{1}}^{\theta\phi}\right)\\ \sqrt{\kappa_{n_{1},m_{1}}^{-1}}\exp\left(j\Phi_{n_{1},m_{1}}^{\phi\theta}\right)&\exp\left(j\Phi_{n_{1},m_{1}}^{\phi\phi}\right)\end{array}\right]\cdot\left[\begin{array}[]{c}F_{\textrm{Tx},s,\theta}\left(\theta_{n_{1},m_{1}}^{\textrm{Tx}},\phi_{n_{1},m_{1}}^{\textrm{Tx}}\right)\\ F_{\textrm{Tx},s,\phi}\left(\theta_{n_{1},m_{1}}^{\textrm{Tx}},\phi_{n_{1},m_{1}}^{\textrm{Tx}}\right)\end{array}\right]$
$\cdot\exp\left(j\frac{2\pi}{\lambda}\left(\hat{r}_{\textrm{Rx},n_{2},m_{2}}\cdot\vec{d}_{\textrm{Rx},u}+\hat{r}_{\textrm{Tx},n_{1},m_{1}}\cdot\vec{d}_{\textrm{Tx},s}\right)\right)\cdot\exp(j2\pi f_{n_{2},m_{2}}t)\cdot\delta{(\tau-\tau_{n_{1},m_{1}}-\tau_{n_{2},m_{2}})}.$
The current main channel modeling method is statistical channel modeling based on GBSM, as mentioned in Section II.B. This method considers the presence of many clusters and multipaths in the wireless channel and describes their delay, angle, power, and other parameters through specific statistical distributions. Given the potential use of RISs in various frequency bands and service scenarios in future wireless networks, it is necessary to develop an RIS channel model that is suitable for full frequency bands and all scenarios. The GBSM-based statistical channel model has the advantage of scalability and can be easily extended to different communication scenarios. Therefore, it is expected to be an effective approach for modeling RIS channels [205].
In [205], the authors propose a GBSM-based channel modeling framework for the RIS cascaded channel, which takes into account the incident angle dependence of the RIS element reflection coefficient. The model was applied to an RIS-assisted MIMO communication scenario, as shown in Fig. 17. The CIR of the RIS cascaded channel between the $s^{\textrm{th}}$ transmitter and the $u^{\textrm{th}}$ receiver is expressed as (24). Similar to the channel model presented in (6), most of the channel parameters in the RIS cascade channel model are defined similarly. However, the RIS cascaded channel model differs in that it includes two subchannels and the RIS response. The specific parameters of the RIS cascaded channel model are listed below.
  • Two sets of cluster and path parameters are included in the equation. $n_{i},m_{i},i\in\{1,2\}$ indicate the identifiers of the cluster and path in the Tx-RIS and RIS-Rx subchannels.
  • The relevant SSPs encompass parameters such as power $P_{n_{1},m_{1}},P_{n_{2},m_{2}}$, doppler shift $f_{n_{2},m_{2}}$, delay $\tau_{n_{i},m_{i}}$, $i\in\{1,2\}$, and angles $\phi_{n_{1},m_{1}}^{\textrm{in}},\theta_{n_{1},m_{1}}^{\textrm{in}},\theta_{n_{1},m_{1}}^{\textrm{Tx}},\phi_{n_{1},m_{1}}^{\textrm{Tx}}$, representing the AOA, EOA, EOD, and AOD, respectively, of the $(n_{1},m_{1})^{\textrm{th}}$ path in the Tx-RIS subchannel. Similarly, $\phi_{n_{2},m_{2}}^{\textrm{Rx}},\theta_{n_{2},m_{2}}^{\textrm{Rx}},\theta_{n_{2},m_{2}}^{\textrm{out}},\phi_{n_{2},m_{2}}^{\textrm{out}}$ denote the AOA, EOA, EOD, and AOD of the $(n_{2},m_{2})^{\textrm{th}}$ path in the RIS-Rx subchannel. $\kappa_{n_{i},m_{i}}$, $i\in\{1,2\}$, represents the XPR of the corresponding path. Additionally, $\Phi_{n_{i},m_{i}}^{p_{1}p_{2}}$, $i\in\{1,2\}$, is the random phase of the $(n_{i},m_{i})^{\textrm{th}}$ path departing in the $p_{1}$ direction and arriving in the $p_{2}$ direction.
  • $F_{\textrm{ris}}^{p_{1}p_{2}}$ represents the RIS overall radiation pattern, with $p_{1},p_{2}\in\{\theta,\phi\}$ denoting vertical and horizontal polarization, respectively. The notation $p_{1}p_{2}$ signifies that the RIS has distinct effects on the cluster or path incident in the $p_{1}$ polarization direction and emitted in the $p_{2}$ polarization direction.
Implementation of the proposed model is based on the 3GPP standardized channel model with some improvements, including three parts: 1) general and large-scale parameters, 2) small-scale parameters, and 3) the CIR for RIS cascaded channel. The model generates large-scale parameters (LSPs) and SSPs for both the Tx-RIS and RIS-Rx subchannels. The overall radiation pattern of the RIS is calculated using a method proposed in [216] to account for the incident angle-dependent characteristics of the RIS reflection coefficient. For further details on the specific implementation processes, readers may refer to [205].
The existing RIS channel modeling approaches can be broadly categorized as statistical approaches [220, 217], deterministic approaches [216, 221], and hybrid approaches [222], in addition to the previously discussed GBSM-based method. These approaches differ in their trade-offs between model accuracy and complexity. Statistical approaches generally have lower model complexity but may sacrifice some accuracy, while deterministic approaches tend to provide higher accuracy but have higher computational complexity. Hybrid models aim to strike a balance between the two, leveraging the advantages of both approaches.
In [220], an RIS-assisted communication channel modeling method for the sub-6 GHz band was proposed based on the GBSM principle, and related work was extended to multiple RIS cases in [217]. In [216], the authors use impedance boundary conditions and equivalent impedance to derive the relationship between the reflection coefficient of the RIS element and the incident angle and then combine the physical optics and RT methods to achieve deterministic channel modeling of the RIS. To validate the accuracy of the proposed model, the authors carried out a path loss measurement activitie in the corridor scenario, and compared the results with RT simulation. The results showed a high level of agreement between the measurement and simulation. Different from the above cluster-based modeling methods, [221] introduced a circuit-based communication model for RIS-assisted wireless systems. This model takes into account the mutual impedance between all radiation elements (transceiver antenna, RIS element) and introduces RIS into the communication model by presenting the influence of RIS on the equivalent current and equivalent voltage of the transceiver antenna. The authors in [222] propose an RIS channel modeling method based on a map-based hybrid channel model (MHCM). In that study, the multipath information in complex scenes was generated by the RT method, and the RIS surface was modeled as a virtual BS excited by different incoming waves. The gain of the RIS for different multipaths was calculated, and finally, the channel coefficients of the entire RIS channel were generated by multipath merging.
VI-D Summary and Prospects
The RIS channel measurement is primarily aimed at measuring the power gain caused by RIS. The characteristics of the RIS channel primarily include reflection coefficient characteristics of the RIS element, RIS radiation pattern characteristics, and RIS cascade loss characteristics. These characteristics are worth considering in RIS channel modeling. Based on these findings, we foresee some challenges that RIS channel modeling research could encounter.
VI-D1 RIS channel measurement
RIS channel measurement plays a crucial role in understanding RIS channel characteristics and validating RIS channel models. However, due to the high transmission loss of the RIS-assisted channel caused by the Tx-RIS-Rx-cascaded link, it imposes greater requirements on the transmission power and dynamic range of the measurement platform. In addition, the measurement platform needs a complete architecture, such as multiple channels, a high sampling rate, and transmit-receive separation, to capture channel characteristics in various complex scenarios.
VI-D2 New characteristics modeling
The reflection coefficient characteristics are important RIS channel characteristics. Various hardware models are previously presented to capture the reflection coefficient characteristic of RIS elements. However, due to the diversity of hardware implementations in RIS elements, it is not feasible to include all types of RIS elements in these hardware models. Considering that it is too expensive to build a hardware model for each type of RIS element, it is a major challenge to propose an accurate and universal RIS hardware model in the future.
VI-D3 RIS channel modeling
The passive characteristic makes RIS have the advantage of low cost and low power consumption. But this also leads to a problem where RIS requires a large area to provide sufficient gain. Therefore, the RIS channel may exhibit near-field effect and SnS like E-MIMO channel. In this case, the RIS response can not be characterized by the RIS overall radiation pattern. Thus, modeling near-field effects and SnS in the RIS channel is a major challenge.
TABLE V: SAGIN measurement campaigns
Scenario Frequency (GHz) Channel Characteristics Reference
Space-to-ground 0.86, 1.55, 39.4 PL [223, 224]
Space-to-ground 0.87, 1.5, 2, 12, 14, 16, 28.2 SF [225, 226, 227, 228, 229, 230, 231]
Space-to-ground 2 Power, SF, capacity, K [232]
Space-to-ground 19, 19.7, 39.04, 39.4 Meteorological effects [233, 234, 235]
Air-to-ground 0.92 PL, DS, Doppler, spatial correlation [236]
Air-to-ground 0.97, 5 PL, DS, K, spatial correlation [237, 238, 239]
Air-to-ground 2.05 PL, DS, excess delay, K [240]
Air-to-ground 2.3 PL [241]
Air-to-ground 4.3 PL, DS, excess delay, SF [242]
Air-to-ground 5 PL, Doppler, K [243]
Air-to-ground 6.5 Power, excess delay, DS, K [244]
Air-to-air 2.4 Power, K [245]
VII SAGIN Channel Measurements and Models
One of the 6G usage scenarios is ubiquitous connectivity by addressing the challenges of connectivity, coverage, capacity, data rate and the mobility of terminals [6]. One focus of this usage scenario is to address presently uncovered or scarcely covered areas, and SAGIN is considered a resolution to achieve worldwide coverage. SAGIN is an integration of satellite systems, aerial networks, and terrestrial communications. Thus in SAGIN, channel characteristics in the communication scenarios of different segments (space, air and ground) should be considered. However, due to the long propagation distance and high node mobility, SAGIN channel characteristics are distinctly different from those of the traditional terrestrial channels [246]. Thus in this section, we review and introduce the present research on channel measurements, characteristics, and models of SAGIN. Since terrestrial systems have been discussed in the above sections, this section is focused more on the SG and AG scenarios. In addition, a brief summary and potential future research directions are provided.
VII-A SAGIN Channel Measurements
For clarity, we classified SAGIN channel measurements, into SG, AG, and air-to-air (AA), and summarized them in Table V.
For SG channels, research work has been started in the early years. Earlier, researchers [223] conducted measurements using the ATS-6 satellite to study the propagation loss of SG links at 860 and 1550 MHz. In [224], propagation loss measurements were carried out at several geographical ground locations in the Q/V frequency band, and a low earth orbit (LEO) satellite channel model was developed using machine learning and statistical methods. A series of measurements of tree attenuation were conducted and studied in various frequency bands such as L-band and K-band [225, 226, 227, 228, 229, 230, 231]. In [232], a satellite MIMO channel study was conducted to investigate the channel capacity in an urban scenario. In [235], meteorological effects, such as rain, clouds, and scintillation, were studied based on field channel measurements in a tropical area. In addition, researchers in [233, 234] conducted measurements to study rain attenuation at the Ka/Q band. According to Table V, it can be found that most of the channel measurements of SG were carried out at L-band, S-band, K-band and Q-band [247], which are the common frequency bands for satellite communications. For channel characteristics, these measurement campaigns mainly focus on the path loss, shadow fading (especially shadow fading caused by tree attenuation), and meteorological effects, and less attention is paid to the small scale characteristics of the channel.
For aerial channels like AG and AA, due to the rise of UAV technology, the corresponding channel measurements have gradually conducted. In [241], the path loss of the UAV channel was studied at 2.3 GHz in urban and nonurban scenarios. In [237, 238, 239], the National Aeronautics and Space Administration (NASA) Glenn Research Center sponsored a comprehensive AG channel measurement study of the C-bands. These measurements covered nearly all the typical local ground environments, including overwater, hilly, mountainous, suburban, and near-urban areas. Researchers in [240] also conducted wideband measurements using a direct-sequence spread-spectrum channel measurement platform to characterize propagation at 2.05 GHz. In [236], a channel measurement for a low-altitude AG channel was conducted at 915 MHz. In [242], AG channel measurements were conducted in open and suburban areas at 4.3 GHz. In [244], authors studied time-varying AG channels based on the measurement data at 6.5 GHz, with a particular focus on the clustering and tracking of multipath components. In [243], the measurement of the AG channel in a near-airport scenario at 5 GHz was conducted. In addition, AA channels for UAVs were investigated based on laboratory measurements and several experimental flights in [245]. According to Table V, it can be found that measurements of aerial channels are conducted mainly in AG scenarios, channel measurements in AA and other scenarios are scarce. Besides, measurement campaigns were conducted in relatively low frequency bands, which are from 0.92 to 6.5 GHz. For channel characteristics, in addition to path loss and shadow fading, AG channel measurements also focus on small scale characteristics such as delay spread, Ricean K-factor (K-factor), Doppler spread and others.
VII-B SAGIN Channel Characteristics
For SAGIN scenarios, especially for SG and AG channels, propagation loss is one of the most important channel characteristics due to the long communication distance. Shadow fading, which is caused by trees or buildings on the ground, could introduce signal attenuation on the order of tens of dB in the received signal. Small-scale characteristics, such as delay spread and Doppler spread, have attracted considerably more attention, especially for LEO satellite wideband communications. In addition, meteorological effects, e.g., rain, snow and fog attenuation, have a greater impact on AG or SG channels compared with terrestrial communication systems. A detailed analysis of the above channel characteristics is given below.
VII-B1 Propagation loss and shadow fading
Because of the long propagation distance in the SG channel, the basic propagation loss of the SG is considered free space propagation, which can be described by the Friis equation [248]. The signal will suffer a large propagation loss of several hundreds of dB. In SG channels, shadow fading is mainly caused by landforms or large objects on the ground, like buildings and trees. The analysis in [227] demonstrated that tree shadowing is a key factor dictating the fade margin for land-mobile satellite systems compared with fading caused by multipath fading. Many studies on shadow fading in SG channels focus on tree shadowing. For example, in [225, 226], the authors found that a single tree led to an additional attenuation ranging from 10 to 20 dB with an average median attenuation of 12 dB. In [228], the additional attenuation values caused by trees with and without leaves were 23 and 7 dB, respectively.
VII-B2 Small-scale characteristics
Small-scale characteristics, including delay spread, Doppler spread, and the K-factor are important for AG channels. In [237], research on AG overwater channels was conducted. The authors found that the delay spread of the AG channel in an overwater scenario is typically small, approximately 10 ns, because the propagation channel is dominated by the LOS path and the reflected path from the water surface. Researchers in [238] found that in a hilly terrain suburban environment, the values of delay spread are mainly within the range of 10-20 ns, but a larger value of nearly 1000 ns also occasionally occurs. For Doppler effect, researchers in [236] conducted channel measurements at 915 MHz and found that a max Doppler shift of 100 MHz at 915 MHz occurs when the UAV is moving directly away from the land station [236]. In [239], the K-factor was measured in urban areas for the L and C-band. The mean value of the K-factor was 12 dB in the L-band and 27.4 dB in the C-band. In addition, the correlation relationships between the K-factor and elevation angle, as well as the K-factor and link distance were investigated in [240] and [243].
VII-B3 Meteorological effects
Compared with terrestrial communication, SG channels could suffer significant meteorological effects, such as atmospheric absorption, rain, fog, and ionospheric effects. These effects could change with frequency, elevation angle, atmospheric pressure, temperature, etc. In [235], the authors conducted experimental channel measurements in the Ku-band in a tropical area to study the impact of various meteorological effects, such as rain, clouds, and tropospheric scintillation. According to the measurement data, the rain attenuation reached approximately 2 dB when the accumulated rainfall (AR) was less than 0.2 mm. When the AR was 0.6 mm, the normalized mean and maximum signal fade levels were 5 and 6.4 dB, respectively.
A series of models have been suggested to predict meteorological effects by ITU-R. In ITU-R P.676 [249], an atmospheric attenuation model is provided, which primarily considers the absorption effects of oxygen and water vapor molecules below 350 GHz. Additionally, ITU-R P.618 offers a method for predicting rain attenuation, which is non-negligible in the mmWave frequency band [250]. In addition, ionospheric effects lead to rapid fluctuations in the amplitude and phase, changes in the polarization angle, and a spread of the delay of the received signal. Models for ionospheric effects, such as ionospheric scintillation and Faraday rotation, are provided in ITU-R P.531 [251].
VII-C SAGIN Channel Modeling
As mentioned in Section II.B, the current main channel modeling method is statistical channel modeling based on GBSM. In the early years, the SAGIN channels, especially SG channels, were modeled by statistically modeling the envelope of the received signal. Generally, the SG channel model can be expressed as the sum of a log-normally distributed random phasor and a Rayleigh phasor [252]:
$r=ze^{j\phi_{0}}+we^{j\phi},$ (25)
where $r$ is the received signal envelope, $ze^{j\phi_{0}}$ represents the LOS component, $we^{j\phi}$ represents other multipaths, the phases $\phi_{0}$ and $\phi$ are uniformly distributed between 00 and $2\pi$, $z$ is log-normally distributed , and $w$ has a Rayleigh distribution. In recent years, meteorological effects have been gradually taken into consideration. In [253], a channel model taking into account weather impairments was proposed. In this case, the PDF of the received signal amplitude is
$p(r)=p_{\beta}(r)p_{w}(r),$ (26)
where $p_{\beta}(r)$ and $p_{w}(r)$ are the PDFs of the mobile fading and weather impairments, respectively. Besides, [235] proposes a tropical weather-aware land-mobile satellite channel model. In this model, multipath fading of SG channels has been modeled based on the Rayleigh distribution, and meteorological effects such as rain, clouds, and tropospheric scintillation are modeled by adjusting the K-factor
$K_{\rm total}=K_{\rm LOS}-K_{\rm Rain}-K_{\rm Cloud}$ (27)
where $K_{\rm LOS}$ is the K-factor under a clear LOS environment, $K_{\rm Rain}$ is the K-factor caused by the rain, and $K_{\rm Cloud}$ is the K-factor caused by the cloud and fog in the environment.
In addition to the above models, 3GPP TR 38.811 [254] provides channel models for non-terrestrial network (NTN) communication (NTN model). The framework of the NTN channel model is similar to [43] with a few modifications. First, the supported scenarios (dense urban, urban, suburban, and rural) and frequency bands (S-band and Ka-band) of the NTN model are modified. Second, the Earth-centered, Earth-fixed (ECEF) coordinate system is used in the NTN model to define the positions of satellites and ground stations instead of the traditional Cartesian coordinate system. Third, the NTN model introduces a circular aperture antenna model as a supplementary antenna configuration. Fourth, the elevation angle of the ground station becomes an important parameter for calculating large-scale and small-scale parameters. Fifth, a series of models are introduced to the NTN model to support the simulation of NTN channel characteristics, such as the clutter loss model used to capture interference from nearby buildings and ground objects. Besides, meteorological effects like atmospheric, rain, and ionospheric fading described previously are supported [250, 249, 251] in the NTN model.
VII-D Summary and Prospects
As discussed above, SAGIN is considered a resolution to achieve the ubiquitous connectivity of 6G, and much research on the SAGIN channel has been carried out. Channel measurements of SAGIN segments like SG and AG have been conducted in many scenarios and frequency bands. Channel characteristics, such as propagation loss, shadow fading, meteorological effects, delay spread, K-factor, Doppler spread, and others have been studied and modeled. However, to improve the capability of SAGIN, further research on the channel is needed.
VII-D1 More SAGIN channel measurements
As discussed above, SAGIN includes satellite systems, aerial networks, and terrestrial communications. A number of channel measurements have been conducted in some segments of SAGIN, such as SG and AG. However, measurements for other segments of SAGIN, such as space-to-space (SS) and space-to-air (SA), are scarce. In addition, SAGIN segments always have large spatial scales. Thus the development of channel measurement platforms that have the ability to measure SAGIN channels is increasingly important. Besides, to obtain and analyze the channel characteristics of the integrated network, channel measurements in scenarios where different segments of SAGIN work together are needed.
VII-D2 Futher exploring for SAGIN channel characteristics
Channel characteristics, such as propagation loss and shadow fading, delay spread, Doppler spread, K-factor, and meteorological effects have been studied in recent years. However, research on other channel characteristics, such as angular spread, is scarce. In addition, because of the lack of measurements, the channel characteristics of some segments, such as SA, SS, and others, have not been deeply studied. Besides, the interaction between different segments of SAGIN should be considered and studied in future research.
VII-D3 SAGIN channel modeling
Channel models for SAGIN segments like SG and AG have been proposed in recent years. For large-scale channel characteristics like propagation loss, channel models of SG and AG can be obtained by modifying the existing propagation loss model. For small-scale channel characteristics, some statistical channel models were proposed in the early years, and 3GPP proposed channel models for simulating fast fading [254]. Besides, ITU provided channel models to capture meteorological effects like atmospheric attenuation, rain attenuation, and others [250, 249, 251]. However, channel models for other segments like SS and SA are still under research. Furthermore, there is a lack of a unified channel model for the integrated SAGIN segments, rather than channel models for a single segment.
VIII 6G Channel Simulator
Channel simulator is the software to generate CIR in different scenarios and configurations based on the existing modeling framework and configuration parameters. Channel simulator plays an important role in designing, evaluating, and optimizing communication systems. In this section, a 6G channel simulator named Beijing University of Posts and Telecommunication and China Mobile Communications Corporation Channel Model Generator-IMT2030 (BUPTCMCCCMG-IMT2030) is introduced. BUPTCMCCCMG-IMT2030 was developed based on the framework of the 5G GBSM standard channel model, and a series of extensions for the 6G technologies and scenarios described in Section IV-VII were integrated. The software structure, main functions, and a few simulation examples including configurations and results of the BUPTCMCCCMG-IMT2030 are given below.
VIII-A Development of Channel Simulators for 6G
In each generation, a few channel simulators were developed and played important roles in performance evaluations of communication technologies. In the 4G era, WINNER II provided the WINNER II channel models and corresponding implementations for link and system-level simulations [66], including generic and clustered delay line models defined for different propagation scenarios. In the 5G era, several organizations have developed simulators for link and system-level channel simulations. In 2014, the Quasi Deterministic Radio Channel Generator (QuaDriGa) was released by Fraunhofer Heinrich-Hertz-Institute [255]. In 2017, New York University (NYU) released a channel simulator named NYUSIM [256] based on the statistical spatial channel model proposed by NYU. In 2018, IMT-2020 CM-BUPT [257] was launched as an implementation of the widely recognized 5G standard GBSM channel model defined in ITU-R M.2412 [42] and 3GPP TR38.901 [43]. In recent years, with the development of 6G research, a series of channel simulators were released to capture channel characteristics of 6G technologies. In 2020, SimRIS was released [258], providing a propagation channel for RIS-assisted communication research in various indoor and outdoor environments. In 2021, researchers from the Higher Institute for Applied Sciences and Technology (HIAST) published the TeraMIMO channel simulator for wideband MIMO THz communications [259]. Besides, in 2023, the NYUSIM was updated to the 4.0 version which supports simulations of THz frequency bands.
Figure 21: Simulation framework for the standard GBSM channel model.
Figure 22: Simulation framework of BUPTCMCCCMG-IMT2030. The blue parts in Module II are simulation steps related to the basic simulation, and the green parts are related to the modifications for 6G.
To meet the requirements of channel simulations supporting higher frequency bands, larger-scale antenna array, ISAC, RIS, and SAGIN, our team developed a 6G channel simulator called BUPTCMCCCMG-IMT2030 based on the framework of the 5G standard GBSM channel model. It supports the generation of CIRs for various scenarios, including urban macrocell (UMa), UMi, rural macrocell (RMa), InH, and SAGIN. New channel features and parameters are also integrated into the platform to support 6G key technologies, such as THz, E-MIMO, ISAC, and RIS.
VIII-B Basic Simulation Framework of BUPTCMCCCMG-IMT2030
BUPTCMCCCMG-IMT2030 is implemented based on the procedure-oriented framework of the 5G standard GBSM channel model with a series of extensions for 6G features. As shown in Fig. 21, the framework of the basic simulation procedure is briefly summarized as follows[43].
TABLE VI: Detail simulation procedure for the astandard GBSM channel model
Steps Descriptions
Step 1 Set the scenario, network layout, and antenna parameters.
Step 2 Assign the propagation condition (LOS/NLOS) according to the network layout and LOS possibility.
Step 3 Calculate the path loss. The path loss models for the calculation are different for different scenarios, frequencies, and propagation conditions.
Step 4 Calculate the LSP, such as DS, AS, K-factor, and shadow fading. The parameters needed for the calculation are correlated to the scenario, frequency, and propagation condition.
Step 5 Generate cluster delays using DS in Step 4.
Step 6 Generate cluster powers based on delays in Step 5.
Step 7–8 Generate arrival and departure angles for clusters and rays based on the AS obtained in Step 4.
Step 9 Generate the XPRs.
Step 10 Draw random initial phases of each ray.
Step 11 Generate the channel coefficient.
Step 12 Apply path loss and shadowing.
VIII-C Software Structure and Functions
A software platform structure of BUPTCMCCCMG-IMT2030 is illustrated in Fig. 22. The software consists of three modules: configuration, simulation, and analysis. In Module I, the configuration module, system parameters, such as scenario, center frequency, and bandwidth, is defined. In Module II, the simulation module, BUPTCMCCCMG-IMT2030, performs a simulation and generates a series of channel parameters. In Module III, the analysis module, simulation results are analyzed and shown in numerical and graphical formats. A stepwise description of the simulation is given as follows.
VIII-C1 Module I–Configuration
In the configuration module, there are two types of configurations: general configurations and 6G feature configurations. General configurations include basic system parameters for a complete simulation, including the scenario, center frequency, bandwidth, positions of BS and UE, antenna type and layout, and mobility of UE. 6G feature configurations are a series of parameters for new channel features for 6G. Various 6G channel features are supported, including an extension of fast-fading parameters and sparsity characteristics for the THz channel, near-field and SnS characteristics for the E-MIMO channel, and other features for channels, i.e., SAGIN, ISAC, and RIS.
TABLE VII: Modifications to the standard GBSM model in BUPTCMCCCMG-IMT2030
6G Features
Steps
Descriptions
THz
Step 3,
Step 4,
Step 11
Step 3: Extend the path loss model for the THz frequency band; Step 4: Extend parameters for the THz frequency band; Step 11: Apply the sparsity to channel coefficients.
E-MIMO
(Near-field
effect, SnS)
Step 1,
Step 11
Step 1: The size of the station region for the antenna array is configured; Step 11: Apply the Near-field effect and SnS to channel coefficients.
ISAC
(Shared cluster,
target scattering)
Step 1,
Step 3-5,
Step 7-12
Step 1: Set the target locations and geometric parameters; Step 3: Add the path loss calculation of the sensing channels; Step 4: Add the correlation between the sensing and communication channels; Step 5: Add the calculation of delays for the sensing channels; Step 7-8: Add the calculation of ray angles, RCS, and clutter for the sensing channels; Step 9-12: Add the calculation of the sensing CIR.
RIS
(Cascade channel)
Step 1-11 Step 1: Set the configuration of RIS; Step 2-10: Propagation condition, path loss, LSP, and SSP are generated for both                  Tx-RIS and RIS-Rx links; Step 11: The cascade channel coefficient of Tx-RIS-Rx is calculated considering the               radiation pattern of RIS.
SAGIN
(SG channel)
Step 1-4 Step 1: Set the configuration of BS and UE in the ECEF coordinate system; Step 2-4: Models and parameters for the SAGIN scenarios are used.
VIII-C2 Module II–Simulation
The simulation module is divided into 2 segments: the basic simulation and modifications for 6G. The basic simulation framework follows that of the 5G standard GBSM channel model, as shown in Fig. 21. The realization of 6G features includes the modifications and extensions of the simulation steps.
The basic simulation segment divides the simulation framework in Fig. 21 into four parts. Part 1 includes steps 1 and 2, which generate general parameters, such as the network layout and link state. Part 2 includes steps 3 and 4, which generate the path loss and large-scale distribution parameters. Part 3 includes steps 5-9, which are used to generate small-scale parameters, such as cluster delays, powers, and angles. Part 4 includes steps 10-12, which generate channel coefficients.
The 6G extensions segment includes five modifications for the 6G new channel features to the basic simulation segment. A summary of the modified steps and descriptions is shown in Table VII.
  • Modifications for the THz channel simulation.
    For the THz channel simulation, steps 3, 4, and 11 are modified. In step 3, the path loss model is extended to the THz frequency band. In step 4, the parameters for generating multipaths are extended to the THz frequency band. In step 11, the model described in [149] is applied to introduce sparsity into the THz channel.
  • Modifications for the E-MIMO channel simulation.
    For the E-MIMO channel simulation, steps 1 and 11 are modified. In step 1, the size of the station region, which is the number of antenna elements within the station region, is configured. In step 11, the spherical wave and the SnS characteristic are applied to calculate the channel coefficient. The phase evolution of the spherical wave is calculated based on the distance between antenna elements and scatterers, which is estimated according to the angles and delays of clusters. The SnS characteristic is simulated by applying the visibility variation of clusters based on a two-state Markov chain.
  • Modifications for the ISAC channel simulation.
    For the ISAC channel simulation, a series of steps are modified. The original steps are used to generate the communication channels, and the modifications are used to generate sensing channels. In step 1, besides the layout of the BS and UE, the location and geometric parameters of the sensing targets are configured. In step 3, the path loss of the sensing channels is calculated according to the configurations of the sensing targets. In step 4, the correlation between the sensing and communication channels is applied. In step 5, the delays of the sensing channels are calculated. In steps 7-8, a series of SSPs of the sensing channels, such as ray angles, RCS, and clutter, are calculated. In steps 9-12, the channel coefficients of both sensing channels and communication channels are calculated.
  • Modifications for the RIS-assisted channel simulation.
    For RIS-assisted channel simulation, BUPTCMCCCMG-IMT2030 has modified a series of steps based on the 5G standard channel model. In the RIS-assisted channel, the channel is divided into two links, Tx-RIS and RIS-Rx, and parameters are generated for them and calculated together. In step 1, the RIS is configured apart from the layout of the BS and UE. In steps 2-10, the propagation condition (LOS/NLOS), path loss, LSPs, and SSPs are generated for both Tx-RIS and RIS-Rx links. In step 11, the cascade channel coefficients of the RIS-assisted channel are calculated by combining the channel coefficients of the Tx-RIS and RIS-Rx links and applying the radiation pattern of the RIS.
  • Modifications for the SAGIN channel simulation.
    For the SAGIN channel simulation, the channel model defined in 3GPP TR 38.811 [254] is implemented and integrated into BUPTCMCCCMG-IMT2030, including modifications to the coordinate system, models, and parameters. In step 1, the ECEF coordinate system is used to locate the BS and UE. In step 2, the propagation conditions (LOS/NLOS) are assigned using a different method. In step 3, the path loss model is modified as described in VII-C. In step 4, the fast-fading parameters for SAGIN are used, which are related to both the frequency and elevation angle.
TABLE VIII: Configurations for simulation for each 6G case
6G Features System Configurations Channel Feature Configurations
THz
Scenario: UMi-street canyon; LOS status: LOS
Center frequency: 132 GHz; Bandwidth: 1.2 GHz
Position (x,y,z): BS:[0,0,3] (m); UE:[10,10,3] (m)
Antenna: BS: Single antenna; UE: Single antenna
Intracluster K-factor: 17.98 dB
E-MIMO
Scenario: UMi-street canyon; LOS status: LOS
Center frequency: 28 GHz; Bandwidth: 200 MHz
Position (x,y,z): BS:[0,0,3] (m); UE:[10,10,3] (m)
Antenna: BS: 256 elements ULA; UE: Single antenna
Stationary region size: 16 elements
ISAC
Scenario: Indoor-office; LOS status: LOS
Center frequency: 28 GHz; Bandwidth: 200 MHz
Position (x,y,z): BS:[0,0,1.5] (m); UE:[10,0,1.5] (m)
Antenna: BS: Single antenna; UE: Single antenna
Number of sensing targets: 3
Number of sensing clusters: 1
Echo path loss: Free space radar path loss
RCS parameters: Fixed
Method of ISAC: Shared cluster
RIS
Scenario: UMi-street canyon; LOS status: LOS
Center frequency: 28 GHz; Bandwidth: 200 MHz
Position (x,y,z): BS:[0,0,3] (m); UE:[10,10,3] (m)
Antenna: BS: Single antenna; UE: Single antenna
RIS size: 32$\times$32 elements
Codebook: Continuous anomalous reflection
ASA: [$10^{\circ}$, $5^{\circ}$, and $1^{\circ}$]
SAGIN
Scenario: Dense urban; LOS status: LOS
Center frequency: 2 GHz (S-band), 28 GHz (Ka-band)
Antenna: BS: Single antenna; UE: Single antenna
Latitude of BS: 0
Longitude of BS: 0
Height of BS: 600-1500 km
Elevation angle: $30^{\circ}$, $60^{\circ}$
Position of UE: Varies with above configurations
VIII-C3 Module III–Analysis
In Module III, a simple analysis of numerical and graphical results based on the generated channel coefficients is provided, including delay spread, angular spread, the reference signal receive power (RSRP), Gini index (used for sparsity analysis), network layout, and CIR of a single link.
VIII-D Applications
In this section, some application examples of the typical 6G technologies are provided. The main simulation configurations for different 6G cases are listed in Table VIII.
VIII-D1 6G Feature 1–THz Frequency Band
Figure 23: Simulation results of channel CIR at THz frequency band (132 GHz) and mmWave frequency band (28 GHz).
Using the configurations in Table VIII, the CIR results of the generated channel at 132 GHz in the THz band were compared with those at 28 GHz in the mmWave band, as shown in Fig. 23. It can be found that the THz channel has 8 clusters and the mmWave channel has 16 clusters. Fewer multipaths are present in the THz band than in the mmWave band, which indicates that the THz channel has a more sparse structure than the mmWave channel. In addition, the delay spread of the THz channel (4.27 ns),is much lower than that of the mmWave channel (30.94 ns) according to Fig. 23. The multipaths of the THz channel are concentrated in a narrower delay range, which means that multipaths in the THz channel are concentrated in a small time period.
VIII-D2 6G Feature 2–E-MIMO
Figure 24: Simulation results of the absolute cross-correlation coefficient between the reference element and other elements along the antenna array.
Based on the configurations in Table VIII, the absolute cross-correlation coefficient of CIRs between the reference element and other elements in comparison with a channel generated using the standard channel model is shown in Fig. 24. With the increase in spacing between two elements, the cross-correlation coefficient of the standard channel model remains stable, but that of the E-MIMO channel continues to decrease, indicating that the correlation of the two elements becomes weaker. In addition, the appearance or disappearance of clusters and the evolution of the phase also occur. The appearance or disappearance of clusters results in a sudden change in the cross-correlation coefficients, including an increase in the appearance and a decrease in the disappearance. When no cluster appears or disappears, the evolution of the phase results in a change in the cross-correlation coefficients.
VIII-D3 6G Feature 3–ISAC
Figure 25: Simulation results of the SD ($\text{SD}_{\text{s}}$) in the sensing channels under different numbers of shared clusters ($N_{0}$) when total clusters ($N$) are 15.
Using the configurations in Table VIII, we can generate the shared and nonshared clusters of the proposed ISAC channel model in (17). The CDF of the SD metric in the sensing channels, which is the PR of the shared clusters ($N_{0}$) to the total clusters ($N$) in the ISAC channel, under different numbers of shared clusters is shown in Fig. 25. A higher SD indicates a greater potential to realize the mutual auxiliary functions of communication and sensing. As $N_{0}$ increases, $\text{SD}_{\text{s}}$ also increases. By controlling the appropriate $N_{0}$ of the proposed channel model, the desired SD value can be achieved, facilitating flexible simulation implementation of ISAC channels.
VIII-D4 6G Feature 4–RIS
Figure 26: Simulation results of instantaneous SNR at the receiver under different channel angular spread for both ideal and non-ideal modulation models of RIS.
Using the configuration in Table VIII, the impact of channel angular spread on the performance of RIS-assisted communication is illustrated in Fig. 26. With the increase in azimuth angle spread of arrival (ASA) in the Tx-RIS channel, the performance of the RIS cascade communication link will worsen. This is because the greater the angular spread, the lower the energy in the specified incident direction. Additionally, the impact of the angle-dependent characteristics of RIS reflection coefficients on RIS-assisted communication, as mentioned in Section VI-B1, has been considered in the simulation. In the legends, ”Non-ideal modulation” represents the simulation results when considering this angle-dependent characteristic. It can be observed that the non-ideal modulation results in approximately 0.5, 0.7, and 1 dB of attenuation under the three angle spreads.
VIII-D5 6G Feature 5–SAGIN
Figure 27: Simulation results of the path loss of the SG channel under different elevation angle configurations.
Using the configurations in Table VIII, the path loss of SG channels with different channel configurations is illustrated in Fig. 27. The path loss of SG channels is affected by the height of satellites, the elevation angle of the link, and the center frequencies. The increase in satellite height results in an increase in propagation distance, which makes the path loss larger. Furthermore, the path loss increases when a higher frequency band is used. Additionally, the elevation angle of the link can affect the path loss in SG scenarios; when the elevation angles increase, the path loss becomes larger.
VIII-E Summary and Prospects
In this section, a channel model simulator called BUPTCMCCCMG-IMT2030 was introduced, including the design framework, user guidelines, and application examples for 6G potential technologies. To summarize, BUPTCMCCCMG-IMT2030 is a channel model simulator developed for 6G wireless communication by the authors. In the current version, BUPTCMCCCMG-IMT2030 supports simulations for the standardized channel model defined in ITU-R M.2412 [42] and adds a series of extensions for 6G features, such as THz, E-MIMO, ISAC, RIS, and SAGIN. Specifically, BUPTCMCCCMG-IMT2030 supports channel simulation scenarios including UMa, UMi, RMa, InH, and SG in SAGIN, and also supports frequency bands including centimeter wave, mmWave, and THz at 132 GHz. In addition, new channel characteristics, such as the SnS of the multiantenna channel and the sparsity of the THz channel, are also included. Simulation frameworks for RIS-assisted and ISAC channels are also implemented. At present, the 2.0 version of BUPTCMCCCMG-IMT2030 has been released on [260]. In future versions, more scenarios (especially more space and air scenarios for SAGIN), frequency bands (more frequency points for THz), and channel characteristics of potential 6G technologies, will be added based on channel measurement and modeling studies to make BUPTCMCCCMG-IMT2030 a powerful assistance tool for channel model simulation.
IX Conclusions
ITU’s current work on the framework and overall objectives represents the official launch of 6G standardization. The technical evaluation work in the standardization is scheduled for 2024, where the 6G channel research is necessary and critical. A survey of 6G channel measurements and modeling and a tutorial of the 6G channel simulator BUPTCMCCCMG-IMT2030 are presented. The representative requirements and challenges of 6G channel research are summarized first. In the survey, we discussed channel measurements and modeling of potential key technologies, including THz, E-MIMO, ISAC, RIS, and SAGIN. Firstly, for channel measurement, a measurement platform with high accuracy and supporting multiple scenarios is urgently needed. Secondly, for channel modeling, new channel characteristics must be properly considered. Geometry-based stochastic modeling is one of the main modeling methods of 6G, but the model needs to be adjusted under different technologies and scenarios. Finally, in the tutorial, we guided the ITU/3GPP GBSM-based 6G channel simulator, BUPTCMCCCMG-IMT2030, which supports the simulation of the aforementioned various 6G potential technologies. The channel simulator has good backward compatibility and is expected to include more scenarios, frequency bands, and channel characteristics in the future.
Acknowledgments
This research is supported in part by National Science Fund for Distinguished Young Scholars (61925102), National Natural Science Foundation of China (92167202, 62101069, 62201086, 62201087), and Beijing University of Posts and Telecommunications-China Mobile Research Institute Joint Innovation Center.